### v3

# Embeddability of joinpowers

Abstract
A general position map $f:K\to M$ of a $k$-dimensional simplicial complex to a $2k$-dimensional manifold (for $k=1$, of a graph to a surface) is a ${\mathbb{Z}}_{2}$-embedding if $|f\sigma\cap f\tau|$ is even for any vertex-disjoint $k$-dimensional faces (for $k=1$, edges) $\sigma,\tau$. We present criteria for ${\mathbb{Z}}_{2}$-embeddability of certain $k$-dimensional complex (for $k=1$, of any graph) to a $2k$-dimensional manifold. These criteria are
$\bullet$ a ‘Kuratowski-type’ version of the Fulek-Kynčl-Bikeev criteria (for $k=1$), and
$\bullet$ a converse to the Dzhenzher-Skopenkov necessary condition (for $k>1$).
Our higher-dimensional criterion allows us to reduce the modulo 2 Kühnel problem on embeddings to a purely algebraic problem.
Our proof is interplay between geometric topology, combinatorics and linear algebra. It is based on calculation of generators in the homology of certain configuration space (the deleted product) of certain complex (joinpower).
MSC 2020: 57Q35, 55N91, 05E45.
Keywords: embedding, mod 2 embedding, joinpower, deleted product, homology.
1 Introduction and main results
Description of main results
We abbreviate ‘$k$-dimensional’ to just ‘$k$-’.
In this paper a complex is the union of closed faces of some simplex. A $k$-complex is the union of some faces of dimension at most $k$ (for $k=1$ this is a graph).
A general position map $K\to M$ of a $k$-complex to a $2k$-manifold is called a ${\mathbb{Z}}_{2}$-embedding if $|f\sigma\cap f\tau|$ is even for any vertex-disjoint $k$-faces $\sigma,\tau$.
We present criteria for ${\mathbb{Z}}_{2}$-embeddability of certain $k$-complex (for $k=1$, of any graph) to $2k$-manifolds.
These criteria are
$\bullet$ a ‘Kuratowski-type’ version of the Fulek-Kynčl-Bikeev criteria, i.e., a version involving only obstructions coming from $K_{5}$ and $K_{3,3}$ (for $k=1$; Theorems 1.1, 1.2 and 1.3), and
$\bullet$ a converse to the Dzhenzher-Skopenkov necessary condition (for $k>1$; Theorem 1.5).
Our higher-dimensional criterion allows us to reduce the modulo 2 Kühnel problem on embeddings to a purely algebraic problem (Conjecture 1.7).
Our proof is an interplay between geometric topology, combinatorics, and linear algebra. It is based on the calculation of generators in the homology of a certain configuration space (the deleted product) of a certain complex (joinpower), see Theorem 3.1, Lemma 3.3 and Remark 1.6.
Some motivation
This subsection is formally not used below.
The classical Heawood inequality gives a lower estimation for a number $g$ such that the complete graph $K_{n}$ on $n$ vertices embeds into the sphere with $g$ handles. A higher-dimensional analogue of the Heawood inequality is the Kühnel conjecture on embeddings. In a simplified form it asks for an analogous estimation of the ‘complexity’ (i.e., of $\dim H_{k}(M;{\mathbb{Z}}_{2})$) of a $2k$-manifold $M$ into which the union of all $k$-faces of $n$-simplex is embedded. See the detailed discussion of the conjecture in [DS22, Remark 1.3], cf. [Ku23, §4].
Recent results on the conjecture work not only for embeddings but for ${\mathbb{Z}}_{2}$-embeddings ([PT19, Theorem 1], [DS22], see [KS21] for a simplified direct proof of [PT19, Theorem 1]). These results involve a necessary condition for ${\mathbb{Z}}_{2}$-embeddability of certain $k$-complexes to $2k$-manifolds (the ‘only if’ parts of Theorems 1.11.3 and 1.5; the ‘only if’ part of Theorem 1.3 is proved similarly but was not stated). This condition is stated (implicitly in [PT19] and explicitly in [DS22, KS21]) in terms of finding the minimal rank of a matrix with certain relations on its entries, and having a certain special form. So this condition is related to the low rank matrix completion problem, and to the Netflix problem from machine learning. Our main result is the sufficiency of this condition (the ‘if’ parts of Theorems 1.11.3 and 1.5). See surveys [DGN+, Sk22n, DS22p], and the references therein.
Known results on embeddability of $k$-complexes in $2k$-manifolds are described in [Sk24, Remark 1.1.7.bc]. Some motivations for studies of ${\mathbb{Z}}_{2}$-embeddability are presented in [KS21e, Remark 1.1.1.a]. The particular case of ${\mathbb{Z}}_{2}$-embeddability of graphs to surfaces is actively studied, see survey [Sc13] and recent papers [Ky16, FK19, Bi21].
For results on configuration spaces of graphs related to Theorem 3.1 see [FH10], [KP11, Theorem 3.16], [Ky16], [AK20, Question (4) from §1.1], and references listed in [Kn18, §1.8, Configuration spaces of graphs]; see [Di08, §19] for higher-dimensional complexes.
Notation and conventions
Below all matrices have ${\mathbb{Z}}_{2}$-entries. Assume that $n\geq 3$. Let $[n]:=\{1,2,\ldots,n\}$.
Denote by $K_{n,n}$ the complete bipartite graph with parts $[n]$ and $[n]^{\prime}:=\{k^{\prime}\ :\ k\in[n]\}$.
Denote by $I_{n}$ the identity matrix of size $n$. Denote by $H_{n}$ the square matrix of size $2n$ whose off-diagonal $2\times 2$-blocks are zero matrices, and whose diagonal $2\times 2$-blocks are $\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$. For a matrix $A$ let $A_{P,Q}$ be the entry in the row of $P$ and column of $Q$.
A (simplicial) $1$-cycle (modulo 2) in a graph $K$ is a set $C$ of edges such that every vertex is contained in an even number of edges from $C$. A (simplicial) $k$-cycle (modulo 2) in a complex $K$ is a set $C$ of $k$-faces such that every $(k-1)$-face is contained in an even number of faces from $C$. We omit ${\mathbb{Z}}_{2}$-coefficients of homology groups. Denote by $H_{k}(K)$ the group of all $k$-cycles in a $k$-complex $K$, with the sum modulo 2 operation. For a $2k$-manifold $M$ denote by $H_{k}(M)$ the modulo 2 homology group, and by $\cap_{M}:H_{k}(M)\times H_{k}(M)\to{\mathbb{Z}}_{2}$ the modulo 2 intersection form (for definitions accessible to non-specialists see [IF, §2], [Sk20, §6, §10]).
Results for graphs
Our results for graphs are Theorems 1.1, 1.2 and 1.3 (of which Theorem 1.1 is a particular case of Theorem 1.3, and Theorem 1.2 is a particular case of both Theorems 1.3 and 1.5).
Let $M_{m}$ be the sphere with $m$ Möbius films, $S_{g}$ the sphere with $g$ handles, and $N$ either $M_{m}$ or $S_{g}$. Denote
$\Omega_{M_{m}}:=I_{m},\quad\beta(M_{m}):=m,\quad\Omega_{S_{g}}:=H_{g},\quad\text{and}\quad\beta(S_{g}):=2g.$
Let $A$ be a symmetric square matrix of $\binom{n}{3}$ rows (whose rows and whose columns correspond to all 3-element subsets of $[n]$). The matrix $A$ is said to be
$\bullet$ independent if $A_{P,Q}=0$ whenever $P\cap Q=\varnothing$;
$\bullet$ additive if
$A_{F-a,Q}+A_{F-b,Q}+A_{F-c,Q}+A_{F-d,Q}=0$
for any 4-element subset $F=\{a,b,c,d\}\subset[n]$ and 3-element subset $Q\subset[n]$ (here we abbreviate $\{a\}$ to $a$).
$\bullet$ non-trivial if for any $5$-element subset $F\subset[n]$ and444 Here and below ‘for any vertex $v$ / edge $e$’ is equivalent to ‘for some vertex $v$ / edge $e$[PT19, Lemma 20], [KS21, Proposition 2.2 and Remark 2.3], [DS22, Remark 4.4].
Strictly speaking, the properties of matrix $A$ should be called ‘$K_{5}$-independence’, ‘$K_{3,3}$-independence’, etc. For the sake of readability, we shorten the notation.
$v\in F$ we have $S_{F,v}A=1$, where $S_{F,v}A$ is the sum of $A_{P,Q}$ over all the $3$ unordered pairs $\{P,Q\}$ of $3$-element subsets $P,Q\subset F$ such that $P\cap Q=\{v\}$.
Theorem 1.1.
A ${\mathbb{Z}}_{2}$-embedding $K_{n}\to N$ exists if and only if there is a $\beta(N)\times{n\choose 3}$-matrix $Y$ such that $Y^{T}\Omega_{N}Y$ is an independent additive non-trivial matrix.
Let $A$ be a symmetric square matrix of $\binom{n}{2}^{2}$ rows (whose rows and whose columns correspond to all cycles of length 4 in $K_{n,n}$). The matrix $A$ is said to be
$\bullet$ independent if $A_{P,Q}=0$ whenever $P$ and $Q$ are vertex-disjoint;
$\bullet$ additive if $A_{P,Q}=A_{X,Q}+A_{Y,Q}$ whenever $P=X\oplus Y$ (as sets $P,X,Y$ of edges).
$\bullet$ non-trivial if for any subgraph $X$ of $K_{n,n}$ isomorphic to $K_{3,3}$ and edge $e$ of $X$ we have $S_{X,e}A=1$, where $S_{X,e}A$ is the sum of $A_{P,Q}$ over all the 2 unordered pairs $\{P,Q\}$ of cycles $P,Q\subset X$ of length 4 whose only common edge is $e$.
Theorem 1.2.
A ${\mathbb{Z}}_{2}$-embedding $K_{n,n}\to N$ exists if and only if there is a $\beta(N)\times{n\choose 2}^{2}$-matrix $Y$ such that $Y^{T}\Omega_{N}Y$ is an independent additive non-trivial matrix.
Theorem 1.3.
A ${\mathbb{Z}}_{2}$-embedding $K\to N$ of a graph $K$ exists if and only if there is a homomorphism $y:H_{1}(K)\to H_{1}(N)$ such that
(independence) $y(P)\cap_{N}y(Q)=0$ whenever $P$ and $Q$ are vertex-disjoint;
($K_{5}$-non-triviality) for any subgraph $F$ homeomorphic to $K_{5}$, and vertex $v$ of $F$ we have $S_{F,v}y=1$, where $S_{F,v}y$ is the sum of $y(P)\cap_{N}y(Q)$ over all the 3 unordered pairs $\{P,Q\}$ of cycles $P,Q\subset F$ of length [^5] Here and in ($K_{3,3}$-non-triviality) the cycles of length 3 and 4 are not the cycles of that length in $K$, but rather they are cycles with 3 and 4, respectively, vertices of degree more than 2 in $F$ and in $X$. 3 such that $v$ is the only common vertex of $P$ and $Q$; and
($K_{3,3}$-non-triviality) for any subgraph $X$ homeomorphic to $K_{3,3}$, and edge $e$ of $X$ we have $S_{X,e}y=1$, where $S_{X,e}y$ is the sum of $y(P)\cap_{N}y(Q)$ over all the 2 unordered pairs $\{P,Q\}$ of cycles $P,Q\subset X$ of length 4 such that $e$ is the only common edge of $P$ and $Q$.
Remark 1.4.
(a) (on the novelty and idea of proof) The ‘only if’ part of Theorem 1.3 is essentially known, and is analogous to [DS22, Remark 1.6], which in turn is analogous to [FK19, Bi21]. Proof of the ‘if’ part is analogous to the proof of the ‘if’ part of Theorem 1.5 for $k=1$ in §4, except that reference to Theorem 3.1.b is replaced by reference to Lemma 3.3. See more comments on the proof in Remark 1.6.c.
(b) In matrix form, the condition of Theorem 1.3 means that there are a set of $r$ generators of $H_{1}(K)$, and a $\beta(N)\times r$-matrix $Y$ such that $A:=Y^{T}\Omega_{N}Y$ is independent additive non-trivial (as defined below).
Let $A$ be a symmetric $r\times r$-matrix whose rows and whose columns correspond to some generators of $H_{1}(K)$. The matrix $A$ is said to be
$\bullet$ independent, if $A_{P,Q}=0$ whenever $P$ and $Q$ are vertex-disjoint;
$\bullet$ additive, if $A_{P,Q}=A_{P_{1},Q}+\ldots+A_{P_{t},Q}$ whenever $P=P_{1}\oplus\ldots\oplus P_{t}$ (as sets of edges).
If $X,Y\in H_{1}(K)$, then express them as a sum of the generators, $X=X_{1}+\ldots+X_{p}$, $Y=Y_{1}+\ldots+Y_{q}$, and define $A_{X,Y}:=\sum_{i,j}A_{X_{i},Y_{j}}$. The matrix $A$ is said to be $K_{5}$-non-trivial ($K_{3,3}$-non-trivial), if the $K_{5}$-non-triviality ($K_{3,3}$-non-triviality) of Theorem 1.3 holds replacing $y(P)\cap_{N}y(Q)$ by $A_{P,Q}$.
The main result for higher dimensions
We consider piecewise-linear (PL) manifolds as defined in [RS72]. Let $M$ be a closed $2k$-manifold. Denote $\beta(M):=\dim H_{k}(M)$. Denote $\Omega_{M}:=I_{\beta(M)}$ if there is $x\in H_{k}(M)$ such that $x\cap_{M}x=1$, and $\Omega_{M}:=H_{\beta(M)/2}$ otherwise, when $\beta(M)$ is known to be even ($\Omega_{M}$ is the matrix of $\cap_{M}$ in some basis).
Denote $D^{j}:=[0,1]^{j}$ and $S^{j-1}:=\partial D^{j}$. A complex $M$ is called $m$-connected if for any $j=0,1,\ldots,m$ every continuous map $f:S^{j}\to M$ extends continuously over $D^{j+1}$.
Let $[n]^{*k+1}$ be the $k$-complex with vertex set $[k+1]\times[n]$, in which every $k+1$ vertices from different lines span a $k$-simplex. For $k=1$ this is $K_{n,n}$. For geometric interpretation see [Ma03, Proposition 4.2.4]. A $k$-octahedron is the set of $k$-faces of a subcomplex (of $[n]^{*k+1}$) isomorphic to $[2]^{*k+1}\cong S^{k}$. For $2$-element subsets $P_{1},\ldots,P_{k+1}\subset[n]$ such a subcomplex is defined by the set $1\times P_{1}\sqcup\ldots\sqcup(k+1)\times P_{k+1}$ of its vertices.
Let $A$ be a symmetric square matrix of $\binom{n}{2}^{k+1}$ rows (whose rows and whose columns correspond to all $k$-octahedra). The matrix $A$ is said to be
$\bullet$ independent if $A_{P,Q}=0$ whenever $P$ and $Q$ are vertex-disjoint;
$\bullet$ additive if $A_{P,Q}=A_{X,Q}+A_{Y,Q}$ whenever $P=X\oplus Y$ (as sets $P,X,Y$ of $k$-faces);
$\bullet$ non-trivial if for any subcomplex $X$ of $[n]^{*k+1}$ isomorphic to $[3]^{*k+1}$ and $k$-face $e$ of $X$ we have $S_{X,e}A=1$, where by $S_{X,e}A$ we denote the sum of $A_{P,Q}$ over all $2^{k}$ unordered pairs $\{P,Q\}$ of $k$-octahedra $P,Q\subset X$ whose only common $k$-face is $e$.
Theorem 1.5.
Let $M$ be a closed $(k-1)$-connected $2k$-manifold. A ${\mathbb{Z}}_{2}$-embedding $[n]^{*k+1}\to M$ exists if and only if there is a $\beta(M)\times{n\choose 2}^{k+1}$-matrix $Y$ such that $Y^{T}\Omega_{M}Y$ is an independent additive non-trivial matrix.
Remark 1.6.
(a) (on the novelty and idea of proof) The ‘only if’ part of Theorem 1.5 is [DS22, Remark 1.9b]. Our main achievement is the ‘if’ part (§4). For this we use a homological ${\mathbb{Z}}_{2}$-embeddability criterion (Theorem 4.3 by E. Kogan; for completeness, we present in §4 a standard deduction from known results). Our main idea is that the independence and non-triviality of the matrix $Y^{T}\Omega_{M}Y$ correspond to generators of the homology group of a certain configuration space (Theorem 3.1).
(b) A version of Theorem 1.5 for the $k$-skeleton of $n$-simplex instead of $[n]^{*k+1}$ is stated analogously (cf. Theorem 1.1). Such a version is presumably false for $k>1$ because the relevant homology group of configuration space is complicated. So Theorem 1.5 is not as trivial a corollary of Theorem 4.3 as it may seem.
(c) Theorem 1.5 generalizes to joins of finite sets of different sizes $n_{1},\ldots,n_{k+1}\geq 3$: a ${\mathbb{Z}}_{2}$-embedding $[n_{1}]*\ldots*[n_{k+1}]\to M$ exists if and only if there is a $\beta(M)\times{n_{1}\choose 2}\ldots{n_{k+1}\choose 2}$-matrix $Y$ such that $Y^{T}\Omega_{M}Y$ is an independent additive non-trivial matrix. The proof of Theorem 1.5 generalizes trivially to a proof of the generalization, see Remarks 2.6, 3.5, 4.5. In those Remarks, the changes of the proofs that repeat changes of the statements are omitted.
Conjecture 1.7.
(a) (asymptotic mod 2 Kühnel conjecture) For any integer $k>0$ there is $c_{k}>0$ such that if $[n]^{*k+1}$ has a ${\mathbb{Z}}_{2}$-embedding to a closed $2k$-manifold $M$, then $\beta(M)\geq c_{k}n^{k+1}$.
By Theorem 1.5, this is equivalent to the following algebraic conjecture:
For any integer $k>0$ there is $c_{k}>0$ such that if there is a $d\times{n\choose 2}^{k+1}$-matrix $Y$ such that either $Y^{T}Y$ or $Y^{T}H_{d/2}Y$ is independent additive non-trivial, then $d\geq c_{k}n^{k+1}$.
Both conjectures are proved for $k=1$ [FK19], and for any $k$ replacing $n^{k+1}$ by $n^{2}$ [DS22].
(b) (analogue of Theorem 1.5 for embeddings) Assume that $k\geq 3$ and $M$ is a closed $(k-1)$-connected $2k$-manifold. An embedding $[n]^{*k+1}\to M$ exists if and only if there is a $\beta(M)\times{n\choose 2}^{k+1}$-matrix $Y$ with integer entries such that $Y^{T}\Omega_{M,{\mathbb{Z}}}Y$ is an independent additive non-trivial matrix.
Here $\Omega_{M,{\mathbb{Z}}}$ is the matrix of the intersection form of $M$ in some basis of $H_{k}(M;{\mathbb{Z}})\cong{\mathbb{Z}}^{\beta(M)}$ (the isomorphism holds since $M$ is $(k-1)$-connected).
Let $A$ be a $(-1)^{k}$-symmetric square matrix with integer entries, of $\binom{n}{2}^{k+1}$ rows. Independence is defined analogously to the mod 2 case. The matrix $A$ is said to be
$\bullet$ additive if $A_{P,Q}=A_{X,Q}+A_{Y,Q}$ whenever $P=X+Y$ (as integer $k$-cycles);
$\bullet$ non-trivial if for any subcomplex $X$ of $[n]^{*k+1}$ isomorphic to $[3]^{*k+1}$ and $k$-face $e$ of $X$ the sum $S_{X,e}A$ is odd.
The ‘only if’ part of this conjecture is proved analogously to [DS22, Theorem 1.8 and Remark 1.9b] for ${\mathbb{Z}}$-embeddings.
2 Generators in homology of the complex
Of this section, only Lemma 2.1 and Theorem 2.3 are used later.
Lemma 2.1.
(a) Any $k$-cycle in $[n]^{*k+1}$ is a sum of some $k$-octahedra.
(b) Any linear relation on $k$-octahedra is a sum of some relations
$K*\{u,v\}+K*\{v,w\}+K*\{w,u\}=0$
for pairwise distinct $u,v,w\in[n]$, and $(k-1)$-octahedra $K$, and relations obtained from that by permutations of ‘join coordinates’ [^6] By * we denote a join. Recall that a $(k-1)$-octahedron is the join $K=K_{1}*\ldots*K_{k}$ of some 2-element subsets $K_{1},\ldots,K_{k}\subset[n]$. The ‘join coordinates’ of $K*\{a,b\}$ are $K_{1},\ldots,K_{k},\{a,b\}$. We take the same permutation of ‘join coordinates’ of $K*\{u,v\}$, of $K*\{v,w\}$, and of $K*\{w,u\}$..
Comment. For $k=1$ this lemma states that
(a) any $1$-cycle in $K_{n,n}$ is a sum of some cycles of length $4$;
(b) any linear relation on such cycles is a sum of some relations $au^{\prime}bv^{\prime}+au^{\prime}bw^{\prime}+av^{\prime}bw^{\prime}=0$ and symmetric to them relations $au^{\prime}bv^{\prime}+au^{\prime}cv^{\prime}+bu^{\prime}cv^{\prime}=0$.
Of this (a) is clear (by Statement 2.4) and is known.
Lemma 2.1 holds by the following ‘dual’ restatement, Lemma 2.2. We use the following analogue of the duality between faces of the regular octahedron and vertices of the cube.
All $k$-faces of the complex $[n]^{*k+1}$ are in 1–1 correspondence (duality) with vertices of the $(k+1)$-cube $[n]^{k+1}$. Then a $k$-octahedron is dual to a parallelepiped defined as a subset $P_{1}\times\ldots\times P_{k+1}$, where $P_{i}$ are 2-element subsets of $[n]$.
coordinate line is a subset of $[n]^{k+1}$ given by fixing all coordinates except one. Two $k$-faces of $[n]^{*k+1}$ have a common face if and only if the two corresponding vertices of $[n]^{k+1}$ belong to one coordinate line. Thus a $k$-cycle in $[n]^{*k+1}$ is dual to a rook cycle defined as a subset of $[n]^{k+1}$ containing an even number of vertices of every coordinate line. The property of being a rook cycle is preserved under modulo $2$ summation of subsets.
Lemma 2.2.
(a) Any rook cycle is a sum of some parallelepipeds.
(b) Any linear relation on parallelepipeds in $[n]^{k+1}$ is a sum of some relations
$P\times\{a,b\}+P\times\{b,c\}+P\times\{c,a\}=0$
for pairwise distinct $a,b,c\in[n]$, and parallelepipeds $P\subset[n]^{k}$, and relations obtained from that by permutations of coordinates.
Proof.
For $a\in[n-1]^{k+1}$ denote by $P(a):=\{n,a_{1}\}\times\ldots\times\{n,a_{k+1}\}$ the parallelepiped with opposite vertices $a$ and $(n,\ldots,n)$.
For (a) it suffices to prove that any rook cycle $C\subset[n]^{k+1}$ equals to the sum $P(C)$ of parallelepipeds $P(a)$ over $a\in C\cap[n-1]^{k+1}$. Let us prove this. The sum $C+P(C)$ is a rook cycle. Since $P(a)\cap[n-1]^{k+1}=\{a\}$, we have $(C+P(C))\cap[n-1]^{k+1}=\varnothing$. Together with any vertex $c$ whose $i$-th coordinate is $n$, the sum $C+P(C)$ contains a vertex $c^{\prime}$ having the same coordinates except the $i$-th coordinate, and having fewer coordinates equal to $n$ than $c$. Hence if non-empty, $C+P(C)$ contains an element from $[n-1]^{k+1}$. Thus $C+P(C)$ is empty, i.e., $C=P(C)$.
Let us prove (b). Any parallelepiped $P$ with some $P_{i}\subset[n-1]$ is a sum of some given relations and two parallelepipeds obtained from $P$ by replacing with $n$ one of the two elements in $P_{i}$. So in each of the two parallelepipeds the number of pairs $P_{j}$ not containing $n$ is smaller than in $P$. Hence any relation between parallelepipeds is a sum of some given relations and a relation $P(a_{1})+\ldots+P(a_{s})=0$ for some pairwise distinct $a_{1},\ldots,a_{s}\in[n-1]^{k+1}$. The latter relation has no summands because
$\varnothing=(P(a_{1})+\ldots+P(a_{s}))\cap[n-1]^{k+1}=\{a_{1},\ldots,a_{s}\}.$
Here the second equality holds because $P(a)\cap[n-1]^{k+1}=\{a\}$. ∎
Let $\widetilde{K_{n}}$ be the graph obtained from $K_{n,n}$ by deleting all ‘diagonal’ edges $jj^{\prime}$, $j\in[n]$. E.g. $\widetilde{K_{3}}$ is the cycle $12^{\prime}31^{\prime}23^{\prime}$ of length $6$.
Let $t:\widetilde{K_{n}}\to\widetilde{K_{n}}$ be the involution switching the parts, i.e., switching $j$ and $j^{\prime}$ for every $j\in[n]$. Denote by $t$ also the induced involution on $H_{1}(\widetilde{K_{n}})$.
[…]
Theorem 2.3.
(a) If $n\geq 4$, then any $1$-cycle in $\widetilde{K_{n}}$ is a sum of some cycles of length $4$.
(b) We have
$\mathop{Ker}(I-t^{\otimes\ell})=\mathop{Im}(I+t^{\otimes\ell})\oplus\left\langle\widetilde{K_{3}}^{\otimes\ell}\right\rangle\subset H_{1}(\widetilde{K_{n}})^{\otimes\ell}.$
For the proof of Theorem 2.3.a we need the following simple (and so known) statement.
Statement 2.4.
Any $1$-cycle in a graph is a sum of some simple cycles for which there are no chords, i.e., edges in the graph between two non-consecutive vertices in the cycle.
Proof.
Any 1-cycle is a sum of some simple cycles. Any chord for a simple cycle of length $l$ gives a representation of the cycle as the sum of two simple cycles of lengths less than $l$. ∎
Proof of Theorem 2.3.a.
A simple cycle in $\widetilde{K_{n}}$ is formed by an $l$-element cyclic sequence of vertices such that
$\bullet$ the number $l$ is even and greater than $2$;
$\bullet$ the parts of vertices alternate in the sequence;
$\bullet$ within each separate part, the vertices are not repeated;
$\bullet$ there are no two consecutive vertices $m$ and $m^{\prime}$.
Assume that such a cycle has no chords. Then any two non-consecutive vertices are not adjacent, i.e., they either are in the same part or are $m$ and $m^{\prime}$. Any vertex has $\frac{l}{2}-2$ non-consecutive vertices from the other part. Then $\frac{l}{2}-2\leq 1$, so $l\in\{4,6\}$.
Figure 1: A cycle of length 6 is the sum of three cycles of length 4
In a simple cycle of length $6$ having no chords, any two opposite vertices are $m$ and $m^{\prime}$. Hence the cycle is $(m_{1},m^{\prime}_{2},m_{3},m^{\prime}_{1},m_{2},m^{\prime}_{3})$ for some pairwise distinct $m_{1},m_{2},m_{3}$. Since $n\geq 4$, there is $a\in[n]-\{m_{1},m_{2},m_{3}\}$. Then the cycle equals
$(m_{1},m_{2}^{\prime},m_{3},a^{\prime})+(m_{2},m_{3}^{\prime},m_{1},a^{\prime})+(m_{3},m_{1}^{\prime},m_{2},a^{\prime}),$
see Figure 1. Now the theorem holds by Statement 2.4. ∎
Lemma 2.5.
Let $S$ be the set of all edges $ij^{\prime}$ of $\widetilde{K_{n}}$ such that $i>j>1$ and $(i,j)\neq(3,2)$. Then the cycles $C_{ij^{\prime}}:=12^{\prime}31^{\prime}ij^{\prime}$, $ij^{\prime}\in S$, form a base of $H_{1}(\widetilde{K_{n}})$.
Proof.
The lemma follows because
$\bullet$ an edge $e^{\prime}\in S$ is contained in the cycle $C_{e}$ if and only if $e=e^{\prime}$;
$\bullet$ the graph $\widetilde{K_{n}}-S$ obtained from $\widetilde{K_{n}}$ by deleting all edges of $S$ is a tree.
The second of these assertions follows because the edges of $\widetilde{K_{n}}-S$ are exactly $32^{\prime}$ and $1i^{\prime},i1^{\prime}$, where $i>1$, so this graph is obtained by connecting with edge $32^{\prime}$ the ‘star’ trees consisting of edges $1i^{\prime}$ with $i>1$, and of edges $1^{\prime}i$ with $i>1$. ∎
Proof of Theorem 2.3.b.
Take the base of $H_{1}(\widetilde{K_{n}})$ given by Lemma 2.5. Then the 1-cycles
$\widetilde{K_{3}}=C_{23^{\prime}},\quad C_{ij^{\prime}},\quad C_{23^{\prime}}+C_{ji^{\prime}},\quad\text{where}\quad i>j>1\quad\text{and}\quad(i,j)\neq(3,2),$
also form a base of $H_{1}(\widetilde{K_{n}})$. Observe that
$C_{23^{\prime}}+C_{ji^{\prime}}=1i^{\prime}j1^{\prime}32^{\prime}12^{\prime}31^{\prime}23^{\prime}=1i^{\prime}j1^{\prime}23^{\prime}=tC_{ij^{\prime}}.$
Thus the group $H_{1}(\widetilde{K_{n}})$ has a base involutively permuted by $t$, with the only fixed vector $\widetilde{K_{3}}$. Hence for any $\ell$, tensor $\ell$-fold products of elements of this base form a base of the group $H_{1}(\widetilde{K_{n}})^{\otimes\ell}$ involutively permuted by $t^{\otimes\ell}$, with the only fixed vector $y_{0}:=\widetilde{K_{3}}^{\otimes\ell}$. Denote this base by
$y_{-m},\ldots,y_{-1},y_{0},y_{1},\ldots,y_{m}\quad\text{so that}\quad t^{\otimes\ell}y_{i}=y_{-i}.$
Thus the subgroups $\mathop{Ker}(I-t^{\otimes\ell})$ and $\mathop{Im}(I+t^{\otimes\ell})$ have bases
$y_{0},y_{1}+y_{-1},\ldots,y_{m}+y_{-m}\quad\text{and}\quad y_{1}+y_{-1},\ldots,y_{m}+y_{-m},$
respectively. This implies Theorem 2.3.b. ∎
Remark 2.6.
Lemmas 2.1 and 2.2 and Theorem 2.3.b generalize to joins of finite sets of different sizes $n_{1},\ldots,n_{k+1}\geq 3$.
The statements of the Lemmas generalize as follows:
$\bullet$ the joinpower $[n]^{*k+1}$ is replaced with $[n_{1}]*\ldots*[n_{k+1}]$,
$\bullet$ the cube $[n]^{k+1}$ is replaced with the parallelepiped $[n_{1}]\times\ldots\times[n_{k+1}]$,
$\bullet$ the numbers $u,v,w,a,b,c$ of parts (b) should belong to $[n_{k+1}]$ not to $[n]$, and in the ’permuted’ relations these numbers should belong to the corresponding $[n_{i}]$.
The proofs of the Lemmas generalize by replacing
$\bullet$ $[n-1]^{k+1}$ with $[n_{1}-1]\times\ldots\times[n_{k+1}-1]$,
$\bullet$ $\{n,a_{1}\}\times\ldots\times\{n,a_{k+1}\}$ with $\{n_{1},a_{1}\}\times\ldots\times\{n_{k+1},a_{k+1}\}$,
$\bullet$ the phrase ’whose $i$-th coordinate is $n$’ with the phrase ’whose $i$-th coordinate is $n_{i}$’,
$\bullet$ the phrase ’having fewer coordinates equal to $n$’ with the phrase ’having $i$-th coordinates equal to $n_{i}$ for fewer $i$’s’,
$\bullet$ $P_{i}\subset[n-1]$ with $P_{i}\subset[n_{i}-1]$.
Both in the statement and in the proof of Theorem 2.3.b the linear space $H_{1}(\widetilde{K_{n}})^{\otimes\ell}$ is replaced with the tensor product of linear spaces $H_{1}(\widetilde{K_{n_{i}}})$, $i\in[\ell]$, and the operator $t^{\otimes\ell}$ is replaced with the tensor product of the corresponding symmetry operators $t_{i}$, $i\in[\ell]$.
3 Generators in homology of the deleted product
Of this section, only Theorem 3.1 and Lemma 3.3 are used in other sections.
In this paper $K$ is a $k$-complex (for $k=1$, a graph). Denote by $E_{k}=E_{k}(K)$ the set of all closed $k$-faces of $K$. Then $E_{k}^{2}$ is the set of all closed $2k$-faces of $K^{2}$.
A (cellular) $2k$-cycle in $K^{2}$ is a subset $C\subset E_{k}^{2}$ such that every $(2k-1)$-face of $K^{2}$ is contained in an even number of $2k$-faces from $C$. In other words, such that for every $(k-1)$-face $a$ and $k$-face $\beta$ of $K$
$\bullet$ there is an even number of $k$-faces $\alpha$ of $K$ such that $\alpha\supset a$ and $(\alpha,\beta)\in C$, and
$\bullet$ there is an even number of $k$-faces $\alpha$ of $K$ such that $\alpha\supset a$ and $(\beta,\alpha)\in C$.
Consider the symmetry (involution) $s$ of $K^{2}$ switching the factors, and the corresponding symmetry of $2k$-cycles in $K^{2}$. A $2k$-cycle is called symmetric if $s$ preserves it.
The combinatorial deleted product of $K$ is
$K^{\times 2}_{\Delta}:=\{(\sigma,\tau)\in E_{k}^{2}\ :\ \sigma\cap\tau=\varnothing\}.$
Theorem 3.1.
(a) The following are symmetric $2k$-cycles in $([n]^{*k+1})^{\times 2}_{\Delta}$:
$\bullet$ ‘symmetrized tori’ $P\times Q+Q\times P$ for vertex-disjoint $k$-octahedra $P,Q$;
$\bullet$ the combinatorial deleted product $([3]^{*k+1})^{\times 2}_{\Delta}$ of the subcomplex $[3]^{*k+1}$.
(b) Any symmetric $2k$-cycle in $([n]^{*k+1})^{\times 2}_{\Delta}$ is a sum of some $2k$-cycles described in (a).
Remark 3.2.
(a) For $k=1$ Theorem 3.1.b states that any symmetric $2$-cycle in $(K_{n,n})^{\times 2}_{\Delta}$ is a sum of some of the following ones: $(K_{3,3})^{\times 2}_{\Delta}$, and ‘symmetrized tori’ $P\times Q+Q\times P$ for vertex-disjoint cycles $P,Q$ of length $4$ in $K_{n,n}$. Cf. [Sa91, §3.4] (observe that the assertion [Sa91, 3.4.2] is wrong, e.g. for disjoint union of two copies of $K_{5}$). See also Lemma 3.3.
(b) Proof of Theorem 3.1.b provides a simpler proof (and a higher-dimensional generalization) of [FH10, the formula for $b_{2}(F(\Gamma,2))$ in Corollary 24]: $\dim H_{2k}(([n]^{*k+1})^{\times 2}_{\Delta})=(n^{2}-3n+1)^{k+1}$ for $n\geq 3$.
Proof of Theorem 3.1 for $k=1$ (formally not used later).
Clearly, any symmetrized torus is a symmetric $2$-cycle. We have
$H_{2}((K_{n,n})^{\times 2}_{\Delta})\overset{(1)}{\cong}H_{2}(\widetilde{K_{n}}\times\widetilde{K_{n}})\overset{(2)}{\cong}H_{1}(\widetilde{K_{n}})^{\otimes 2},\quad\text{where}$
$\bullet$ $H_{2}(X)$ is the group of cellular $2$-cycles in $X$;
$\bullet$ isomorphism (1) is induced by the 1–1 correspondence $\alpha_{1}\alpha_{2}^{\prime}\times\beta_{1}\beta_{2}^{\prime}\mapsto\alpha_{1}\beta_{1}^{\prime}\times\alpha_{2}\beta_{2}^{\prime}$ of $2$-cells, for $\alpha,\beta\in[n]^{2}$ such that $\alpha_{1}\neq\beta_{1}$ and $\alpha_{2}\neq\beta_{2}$; here for $a,b\in[n]$ we denote by $ab^{\prime}$ the edge joining $a$ and $b^{\prime}$; if one of $\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}$ is replaced by empty set, then the formula gives a 1–1 correspondence of edges, so that the two 1–1 correspondences respect incidence; this gives a 1–1 correspondence of 2-cycles(note that there are no 1–1 correspondences between vertices, so this construction gives no 1–1 correspondence of 1-cycles);
$\bullet$ isomorphism (2) is the Künneth formula, and so maps the 2-cycle $y\times z$ to $y\otimes z$.
This composition of isomorphisms maps
$\bullet$ the symmetry on $(K_{n,n})^{\times 2}_{\Delta}\subset K_{n,n}\times K_{n,n}$ to $t\otimes t$; and
$\bullet$ the product of disjoint cycles $A*B$ and $C*D$ of length $4$ in $K_{n,n}$ to $(A*C)\otimes(B*D)$, where $A,B,C,D$ are $2$-element subsets of $[n]$.
The inverse of this composition maps $\widetilde{K_{3}}\otimes\widetilde{K_{3}}$ to $(K_{3,3})^{\times 2}_{\Delta}$, which completes the proof of (a).
The elements of $H_{1}(\widetilde{K_{n}})^{\otimes 2}$ invariant under the involution $t\otimes t$ form the subgroup $\mathop{Ker}(I-t\otimes t)$. By Theorem 2.3.a $\mathop{Im}(I+t\otimes t)$ is generated by the ‘symmetrized tori’ $a\otimes b+t(a)\otimes t(b)\in H_{1}(\widetilde{K_{n}})^{\otimes 2}$, for cycles $a,b$ of length 4 in $\widetilde{K_{n}}$. Thus (b) follows by Theorem 2.3.b for $\ell=2$. ∎
Proof of Theorem 3.1 for arbitrary $k$.
Clearly, any symmetrized torus is a symmetric $2k$-cycle. We have
$H_{2k}(([n]^{*k+1})^{\times 2}_{\Delta})\overset{(1)}{\cong}H_{k+1}(\widetilde{K_{n}}^{\times k+1})\overset{(2)}{\cong}H_{1}(\widetilde{K_{n}})^{\otimes k+1},\quad\text{where}$
$\bullet$ $H_{j}(X)$ is the group of (cellular) $j$-cycles in $X$; cellular $j$-cycles in $K^{j}$ are defined analogously to $j=2$;
$\bullet$ isomorphism (1) is induced by the 1–1 correspondence [^7] Recall that for $2$-element subsets $P_{1},\ldots,P_{k+1}\subset[n]$ the $k$-faces $a_{1}*\ldots*a_{k+1}$, $a_{i}\in P_{i}$, of the $k$-octahedron $P_{1}*\ldots*P_{k+1}$ are spanned by vertices $(i,a_{i})$. $(\alpha_{1}*\ldots*\alpha_{k+1})\times(\beta_{1}*\ldots*\beta_{k+1})\mapsto(\alpha_{1}*\beta_{1})\times\ldots\times(\alpha_{k+1}*\beta_{k+1})$ between $2k$-cells and $(k+1)$-cells for $\alpha,\beta\in[n]^{k+1}$ such that $\alpha_{j}\neq\beta_{j}$ for every $j\in[k+1]$; if one of $\alpha_{1},\ldots,\alpha_{k+1},\beta_{1},\ldots,\beta_{k+1}$ is replaced by empty set, then the formula gives a 1–1 correspondence between $(2k-1)$-cells and $k$-cells, so that the two 1–1 correspondences respect incidence; this gives a 1–1 correspondence between $2k$-cycles and $(k+1)$-cycles;
$\bullet$ isomorphism (2) is the Künneth formula, and so maps the $(k+1)$-cycle $y_{1}\times\ldots\times y_{k+1}$ to $y_{1}\otimes\ldots\otimes y_{k+1}$.
This composition of isomorphisms maps
$\bullet$ the symmetry on $([n]^{*k+1})^{\times 2}_{\Delta}\subset[n]^{*k+1}\times[n]^{*k+1}$ to $t^{\otimes k+1}$; and
$\bullet$ the product of disjoint $k$-octahedra
$A_{1}*\ldots*A_{k+1},D_{1}*\ldots*D_{k+1}\subset[n]^{*k+1}\quad\text{to}\quad(A_{1}*D_{1})\otimes\ldots\otimes(A_{k+1}*D_{k+1}),$
where $A_{1},\ldots,A_{k+1},D_{1},\ldots,D_{k+1}$ are $2$-element subsets of $[n]$.
The inverse of this composition maps $\widetilde{K_{3}}^{\otimes k+1}$ to $([3]^{*k+1}))^{\times 2}_{\Delta}$, which completes the proof of (a).
The elements of $H_{1}(\widetilde{K_{n}})^{\otimes k+1}$ invariant under the involution $t^{\otimes k+1}$ form the subgroup $\mathop{Ker}(I-t^{\otimes k+1})$. By Theorem 2.3.a $\mathop{Im}(I+t^{\otimes k+1})$ is generated by the ‘symmetrized tori’ $a_{1}\otimes\ldots\otimes a_{k+1}+t(a_{1})\otimes\ldots\otimes t(a_{k+1})\in H_{1}(\widetilde{K_{n}})^{\otimes k+1}$, for cycles $a_{1},\ldots,a_{k+1}$ of length $4$ in $\widetilde{K_{n}}$. Thus (b) follows by Theorem 2.3.b for $\ell=k+1$. ∎
Let $K$ be a graph homeomorphic to $K_{5}$ or $K_{3,3}$. For an edge $\sigma$ of $K$ denote by $[\sigma]$ the union of all edges of $K$ contained in the edge of $K_{5}$ or $K_{3,3}$ containing $\sigma$. The economic deleted product of $K$ is
$\{(\sigma,\tau)\in E_{1}^{2}\ :\ [\sigma]\cap[\tau]=\varnothing\}.$
Lemma 3.3 (cf. Remark 3.2.a).
For a graph $K$ any symmetric $2$-cycle in $K^{\times 2}_{\Delta}$ is a sum of some of the following ones:
$\bullet$ $P\times Q+Q\times P$ for vertex-disjoint cycles $P,Q$,
$\bullet$ the economic deleted product of a subgraph homeomorphic to $K_{5}$ or to $K_{3,3}$.
The following lemma is used in the proof of Lemma 3.3. Let $L$ be a (cell) complex and $t$ an involution on $L$. A (cellular) $k$-cycle $C$ in $L$ is said to be
$\bullet$ $t$-symmetric if $t$ preserves $C$;
$\bullet$ $t$-minimal if $C$ is inclusion-minimal among non-zero $t$-symmetric $k$-cycles.
$\bullet$ minimal if $C$ is inclusion-minimal among all (not necessarily $t$-symmetric) non-zero $k$-cycles.
Lemma 3.4.
(a) Any $t$-symmetric $k$-cycle is a sum of $t$-minimal ones.
(b) Any $t$-minimal $k$-cycle equals either $C$ or $C+tC$, where $C$ is a minimal $k$-cycle.
Proof.
(a) Denote the given $t$-symmetric $k$-cycle by $C$. Use induction on the number $|C|$ of cells in $C$.
The base $|C|=0$ is obvious.
Suppose that $|C|>0$ and that (a) holds for every $k$-cycle having fewer cells. If $C$ is $t$-minimal, then (a) holds for $C$. Otherwise, $C$ contains a $t$-symmetric $k$-cycle $C_{1}$ distinct from both 00 and $C$. Also $C$ contains a $t$-symmetric $k$-cycle $C_{2}:=C-C_{1}\neq C$. Thus, $|C_{1}|<|C|$ and $|C_{2}|<|C|$. So (a) holds both for $C_{1}$ and for $C_{2}$. Hence, (a) holds for $C=C_{1}+C_{2}$. ∎
(b) Let $C_{0}$ be a $t$-minimal $k$-cycle.
If $C_{0}$ is minimal, then (b) is clear. Otherwise, $C_{0}$ contains a $k$-cycle $C$ distinct from both 00 and $C_{0}$. Note that $C_{0}=tC_{0}\supset tC$. Hence $C_{0}\supset C+tC$. Since $C_{0}$ is $t$-minimal, $C$ is not $t$-symmetric, so, $C+tC\neq 0$. Thus the $t$-symmetric $k$-cycle $C+tC$ equals $C_{0}$.
It remains to prove that $C$ is minimal. Note that $C\cap tC\subset C\subset C_{0}=C+tC$, so $C\cap tC=\varnothing$. Suppose to the contrary that $C$ contains a $k$-cycle $D$ distinct both from 00 and from $C$. Then $tD\subset tC$,  $D\cap tD\subset C\cap tC=\varnothing$,  $0\neq D+tD\subsetneq C+tC=C_{0}$, and $D+tD$ is $t$-symmetric. A contradiction to the $t$-minimality of $C_{0}$. ∎
Proof of Lemma 3.3.
(This is reduction to [Sa91, 3.4.1]; see a direct proof in §5.)
By Lemma 3.4.a, every symmetric $2$-cycle in $K^{\times 2}_{\Delta}$ is a sum of $s$-minimal $2$-cycles. Let $C_{0}$ be any of the $s$-minimal $2$-cycles. By Lemma 3.4.b, $C_{0}$ is either $C$ or $C+sC$, where $C$ is a minimal $2$-cycle. By [Sa91, 3.4.1],
every minimal $2$-cycle in $K^{\times 2}_{\Delta}$ is either a product of vertex-disjoint cycles, or the economic deleted product of a subgraph homeomorphic to $K_{5}$ or to $K_{3,3}$.
If $C$ is a product of vertex-disjoint cycles $P$ and $Q$, then $C$ is not symmetric, so, $C_{0}=C+sC=P\times Q+Q\times P$.
If $C$ is such a deleted product, then $C$ is symmetric. Hence $C+sC=0\neq C_{0}$. So, $C_{0}=C$ is the same deleted product. ∎
Remark 3.5.
Theorem 3.1 generalizes to joins of finite sets of different sizes $n_{1},\ldots,n_{k+1}\geq 3$.
The statement generalizes by replacing $[n]^{*k+1}$ with $[n_{1}]*\ldots*[n_{k+1}]$.
The proof generalizes by replacing
$\bullet$ $[n]^{k+1}$ with the product of the sets $[n_{i}]$ over $i\in[k+1]$,
$\bullet$ the complex $\widetilde{K_{n}}^{\times k+1}$ with the product of the graphs $\widetilde{K_{n_{i}}}$ over $i\in[k+1]$,
$\bullet$ $H_{1}(\widetilde{K_{n}})^{\otimes k+1}$ with the tensor product of the linear spaces $H_{1}(\widetilde{K_{n_{i}}})$ over $i\in[k+1]$,
$\bullet$ $t^{\otimes k+1}$ with the tensor product of the corresponding symmetry operators $t_{i}$ over $i\in[k+1]$,
$\bullet$ the phrase ’$A_{1},\ldots,A_{k+1},D_{1},\ldots,D_{k+1}$ are $2$-element subsets of $[n]$’ with the phrase ’$A_{i},D_{i}$ are $2$-element subsets of $[n_{i}]$ for any $i\in[k+1]$’.
4 Proof of the ‘if’ part of Theorem 1.5
In this section $H$ is any linear space over ${\mathbb{Z}}_{2}$, and $\cdot:H\times H\to{\mathbb{Z}}_{2}$ is any symmetric bilinear form.
Lemma 4.1.
Let $Y_{1},\ldots,Y_{p}\in H$ be vectors, and $P_{1},\ldots,P_{p}$ generators of a linear space $V$ over ${\mathbb{Z}}_{2}$. Assume that if a sum of some $P_{j}$ is zero, then the $\cdot$-product of the sum of the corresponding $Y_{j}$, and $Y_{i}$, is zero for any $i\in[p]$. Then there is a homomorphism $\psi:V\to H$ such that $\psi P_{a}\cdot\psi P_{b}=Y_{a}\cdot Y_{b}$ for every $a,b\in[p]$.
Proof.
 [^8] Observe that the map $P_{a}\mapsto Y_{a}$ need not extend to a homomorphism $V\to H$.
We may assume that $P_{1},\ldots,P_{t}$ is a basis of $V$. Define a homomorphism $\psi:V\to H$ by setting $\psi(P_{j}):=Y_{j}$ for every $j\in[t]$.
Express any $P_{a}$ as a sum of some basic elements: $P_{a}=\sum_{j}P_{j}$. Then $\psi P_{a}-Y_{a}=\sum_{j}Y_{j}-Y_{a}$
$\bullet$ is a sum of some of vectors $Y_{1},\ldots,Y_{p}$;
$\bullet$ has zero $\cdot$-product with $Y_{i}$ for any $i\in[p]$ (by the hypothesis).
Hence
$\psi P_{a}\cdot\psi P_{b}-Y_{a}\cdot Y_{b}=(\psi P_{a}-Y_{a})\cdot Y_{b}+Y_{a}\cdot(\psi P_{b}-Y_{b})+(\psi P_{a}-Y_{a})\cdot(\psi P_{b}-Y_{b})=0.$
For a map $y:E_{k}\to H$ define
$\bullet$ a homomorphism $\widehat{y}:H_{k}(K)\to H$ by $\widehat{y}(A)=\sum\limits_{\sigma\in A}y(\sigma)$;
$\bullet$ $y^{2}(C,\cdot):=\sum\limits_{\{\sigma,\tau\}\in C/s}y(\sigma)\cdot y(\tau)\in{\mathbb{Z}}_{2}$, where $C$ is a symmetric $2k$-cycle $C$ in $K^{2}$.
Lemma 4.2.
(a) For any homomorphism $\psi:H_{k}(K)\to H$ there is a map $y:E_{k}\to H$ such that $\psi=\widehat{y}$.
(b) Let $y:E_{k}\to H$ be a map. For $k$-cycles $P_{1},\ldots,P_{q}$, $Q_{1},\ldots,Q_{q}$ in $K$ take a symmetric $2k$-cycle $C_{P,Q}:=\sum\limits_{j=1}^{q}(P_{j}\times Q_{j}+Q_{j}\times P_{j})$ in $K^{2}$. Then $y^{2}(C_{P,Q},\cdot)=\sum\limits_{j=1}^{q}\widehat{y}(P_{j})\cdot\widehat{y}(Q_{j})$.
Part (a) is simple and known. (Part (a) holds because the scalar product $H^{k}(K;H)\times H_{k}(K;H)\to H$ is nondegenerate. In other words, part (a) follows by the Universal Coefficients Theorem because $\psi$ can be regarded as an element of $H^{k}(K;H)$, and this element is such a map $y$ up to cohomology.) Part (b) is proved by direct calculation.
Denote by $|S|_{2}\in{\mathbb{Z}}_{2}$ the number of elements modulo 2 in a finite set $S$.
Let $f:K\to{\mathbb{R}}^{2k}$ be a general position PL map (see definition in [Sk24, Definition 1.1.6.a]). For a symmetric $2k$-cycle $C$ in $K^{\times 2}_{\Delta}$ let the $C$-van Kampen number of $K$ be
$v(C)=v_{K}(C):=\sum\limits_{\{\sigma,\tau\}\in C/s}|f\sigma\cap f\tau|_{2}\in{\mathbb{Z}}_{2},$
Recall that $v_{K}(C)$ depends only on $K,C$, not on $f$, cf. [Sk18, Remark 1.5.5.c].
Theorem 4.3 (E. Kogan; [KS21e, (E)$\Leftrightarrow$(R’) of Proposition 2.5.1]).
Let $M$ be a closed $(k-1)$-connected $2k$-manifold. There is a ${\mathbb{Z}}_{2}$-embedding $K\to M$ if and only if
(R’) there is a map $y:E_{k}\to H_{k}(M)$ such that for any symmetric $2k$-cycle $C$ in $K^{\times 2}_{\Delta}$ we have $v(C)=y^{2}(C,\cap_{M})$.
Theorem 4.3 follows because (as we prove below) (R’) is equivalent to the condition (RH) of Theorem 4.4 below.
Proof of the ‘if’ part of Theorem 1.5.
Take
$V=H_{k}([n]^{*k+1}),\quad H=H_{k}(M),\quad\cdot=\cap_{M},\quad p={n\choose 2}^{k+1},$
and different $k$-octahedra $P_{1},\ldots,P_{p}\in V$. Take a base of $H_{k}(M)$ in which $\cap_{M}$ has matrix $\Omega_{M}$. For any $a\in[p]$ let $Y_{a}\in H_{k}(M)$ be the vector expressed in this base as the column of $Y$ corresponding to $P_{a}$. Then for any $k$-octahedra $P_{a},P_{b}$ we have $Y_{a}\cap_{M}Y_{b}=(Y^{T}\Omega_{M}Y)_{P_{a},P_{b}}$. So by Lemma 2.1 and the additivity of $Y^{T}\Omega_{M}Y$ the assumptions of Lemma 4.1 are fulfilled. We obtain a homomorphism $\psi:H_{k}([n]^{*k+1})\to H_{k}(M)$ such that
$(*)\qquad\psi P\cap_{M}\psi Q=(Y^{T}\Omega_{M}Y)_{P,Q}\quad\text{for any $k$-octahedra}\quad P,Q.$
By Lemma 4.2.a there is a map $y:E_{k}\to H_{k}(M)$ such that $\psi=\widehat{y}$. By Theorems 4.3 and 3.1.b it suffices to check the equality $y^{2}(C,\cap_{M})=v(C)$ only for $2k$-cycles $C$ from Theorem 3.1.a.
Let $C=P\times Q+Q\times P$ for vertex-disjoint $k$-octahedra $P,Q$. Then
$y^{2}(C,\cap_{M})\overset{(1)}{=}\psi(P)\cap_{M}\psi(Q)\overset{(2)}{=}0\overset{(3)}{=}|f(P)\cap f(Q)|_{2}\overset{(4)}{=}v(C),\quad\text{where}$
$\bullet$ (1) holds by Lemma 4.2.b;
$\bullet$ (2) holds by (*) and the independence;
$\bullet$ (3) holds by the Parity Lemma (asserting that every two general position $k$-cycles in ${\mathbb{R}}^{2k}$ intersect at an even number of points, cf. [Sk, §1.3 and §4.8 ‘Algebraic intersection number’]);
$\bullet$ (4) holds by the definition of $v(C)$.
Let $C=([3]^{*k+1})^{\times 2}_{\Delta}$. Then $C=\sum(P\times Q+Q\times P)$ is the sum over all $2^{k}$ unordered pairs $\{P,Q\}$ of $k$-octahedra $P,Q\subset[3]^{*k+1}$ whose only common $k$-face is $1^{*k+1}$ [DS22, Lemma 2.3]. Hence
$y^{2}(C,\cap_{M})\overset{(1)}{=}\sum\psi(P)\cap_{M}\psi(Q)\overset{(2)}{=}S_{[3]^{*k+1},1^{*k+1}}Y^{T}\Omega_{M}Y\overset{(3)}{=}1\overset{(4)}{=}v(C),\quad\text{where}$
$\bullet$ (1) holds by Lemma 4.2.b;
$\bullet$ (2) holds by (*) and definition of $S_{X,e}$;
$\bullet$ (3) holds by the non-triviality; and
$\bullet$ (4) holds by [vK32, Satz 5] (for an alternative modern proof see [DS22, §4, proof of Lemma 4.2]). ∎
Theorem 4.4.
Let $M$ be a closed $(k-1)$-connected $2k$-manifold. There is a ${\mathbb{Z}}_{2}$-embedding $K\to M$ if and only if
(RH) there is a homomorphism $\psi:H_{k}(K)\to H_{k}(M)$ such that $\omega(\psi)=v_{K}$ (see the required definitions below).
This is [PT19, Proposition 21 and Theorem 15], [Sk24, Theorem 1.1.5 and Lemma 2.1.1].
Let $f:K\to{\mathbb{R}}^{2k}$ be a general position PL map. Take any pair of vertex-disjoint $k$-faces $\sigma,\tau$ of $K$. By general position the intersection $f\sigma\cap f\tau$ consists of a finite number of points. Assign to the pair $\{\sigma,\tau\}$ the residue
$\nu(f)\{\sigma,\tau\}:=|f\sigma\cap f\tau|_{2}.$
Denote by $K^{*}:=K^{\times 2}_{\Delta}/s$ the set of all unordered pairs of non-adjacent $k$-faces of $K$.
The obtained map $\nu(f):K^{*}\to{\mathbb{Z}}_{2}$ is called the (modulo 2) intersection cocycle of $f$.
Let $H^{2k}(K^{*})$ be the group of modulo 2 cohomology classes of maps (=cocycles) $K^{*}\to{\mathbb{Z}}_{2}$.
It is classical [vK32] that the class
$v_{K}:=[\nu(f)]\in H^{2k}(K^{*})$
is well-defined, i.e. is independent of the choice of $f$.
Take a map $y:E_{k}\to H_{k}(M)$. Assign to any pair $\{\sigma,\tau\}$ of vertex-disjoint $k$-faces $\sigma,\tau$ of $K$ the residue
$\omega(y)\{\sigma,\tau\}:=y(\sigma)\cap_{M}y(\tau).$
We obtain a map $\omega(y):K^{*}\to{\mathbb{Z}}_{2}$.
By Lemma 4.2.a for any homomorphism $\psi:H_{k}(K)\to H_{k}(M)$ there is a map $y:E_{k}\to H_{k}(M)$ such that $\psi=\widehat{y}$. Define
$\omega(\psi):=[\omega(y)]\in H^{2k}(K^{*}).$
It is easy to check that $\omega(\psi)$ is well-defined, i.e. that maps $\omega(y)$ are cohomologous for two maps $y$ giving the same homomorphism $\psi=\widehat{y}$ (because such maps $y$ are cohomologous).
Denote by $\frown:H^{2k}(K^{*})\times H_{2k}(K^{*})\to{\mathbb{Z}}_{2}$ the ‘scalar’ product.
Proof that (RH)$\Rightarrow$(R’).
Take a general position PL map $f\colon K\to{\mathbb{R}}^{2k}$. By Lemma 4.2.a there is a map $y:E_{k}\to H_{k}(M)$ such that $\psi=\widehat{y}$. Then by (RH)
$y^{2}(C,\cap_{M})=[\omega(y)]\frown C/s=[\nu(f)]\frown C/s=v(C).$
Proof that (R’$\Rightarrow$(RH).
Take a general position PL map $f\colon K\to{\mathbb{R}}^{2k}$. Let $\psi=\widehat{y}$. Then
$\omega(\psi)=[\omega(y)]=[\nu(f)]=v_{K}.$
Here the middle equality holds because the ‘scalar’ product $\frown$ is nondegenerate, and for any symmetric $2k$-cycle $C$ in $K^{\times 2}_{\Delta}$ by (R’) we have
$\sum_{\{\sigma,\tau\}\in C/s}(\omega(y)-\nu(f))\{\sigma,\tau\}=y^{2}(C,\cap_{M})-v(C)=0.$
Remark 4.5.
Theorem 1.5 generalizes to joins of finite sets of different sizes $n_{1},\ldots,n_{k+1}\geq 3$ (see Remark 1.6.c). The proof of the ‘if’ part of Theorem 1.5 generalizes by replacing $[n]^{*k+1}$ with $[n_{1}]*\ldots*[n_{k+1}]$ and ’$p={n\choose 2}^{k+1}$’ with ’$p={n_{1}\choose 2}\ldots{n_{k+1}\choose 2}$’.
5 Appendix: a direct proof of Lemma 3.3
In this section $L$ is a graph.
The support (the projection) of a $2$-cycle $C$ in $L^{2}$ is the subgraph of $L$ consisting of all edges which occur as one of the factors of the 2-cells of $C$.
Lemma 5.1 (cf. [Sa91, p. 86 above]).
Suppose that $L^{\times 2}_{\Delta}$ contains exactly one non-zero symmetric $2$-cycle $C$, and $L$ is the support of $C$. Then
$\bullet$ $L$ is homeomorphic to either $K_{5}$ or $K_{3,3}$ or $P\sqcup Q$ for some vertex-disjoint cycles $P,Q$, and
$\bullet$ the $2$-cycle $C$ is the economic deleted product of $L$ in the first two cases, and is $P\times Q+Q\times P$ in the last one.
Lemma 5.2 (cf. [Sa91, bottom of p. 86]).
If $L^{\times 2}_{\Delta}$ contains an $s$-minimal $2$-cycle with support $L$, then $L^{\times 2}_{\Delta}$ contains exactly one non-zero symmetric $2$-cycle.
Proof of Lemma 3.3.
(This proof is essentially a symmetric version of a well-structured exposition of the proof from [Sa91, §3.3].) By Lemma 3.4.a, every symmetric $2$-cycle in $K^{\times 2}_{\Delta}$ is a sum of $s$-minimal $2$-cycles. Let $C$ be any of the summands. Let $L\subset K$ be the support of $C$. Then $C$ is $s$-minimal in $L$. So by Lemma 5.2, $C$ is the only non-zero symmetric $2$-cycle in $L$. Hence Lemma 3.3 holds by Lemma 5.1. ∎
To prove Lemmas 5.1 and 5.2, the following Lemmas are required.
Lemma 5.3.
If there exists a symmetric $2$-cycle in $L^{\times 2}_{\Delta}$ with support $L$, then in $L$,
(a) any edge is vertex-disjoint with some simple cycle;
(b) any edge lies in some simple cycle;
(c) for any vertex of degree $2$ there exists a simple cycle vertex-disjoint with both edges issuing out of the vertex.
Proof.
Denote given $2$-cycle by $C$. For any edge $e$ let $C_{e}$ be the set of all edges $e^{\prime}$ such that $(e,e^{\prime})\in C$.
Since the support of $C$ is $L$, for any edge $e$ the set $C_{e}$ is a nonempty $1$-cycle vertex-disjoint with $e$, and any edge lies in some $C_{e}$. So, we proved (a) and (b).
In order to prove (c) denote by $e_{1},e_{2}$ the edges issuing out of a vertex of degree 2. For any edge $e$, we have $(e_{1}\in C_{e})\Leftrightarrow(e_{2}\in C_{e})$, i. e., $(e\in C_{e_{1}})\Leftrightarrow(e\in C_{e_{2}})$. Hence $C_{e_{1}}=C_{e_{2}}$ is a nonempty $1$-cycle vertex-disjoint both with $e_{1}$ and with $e_{2}$. ∎
A graph obtained from a cycle graph $(V,E)$ by adding a new vertex $v_{0}$ and edges $v_{0}v$ for all $v\in V$ is called a wheel.
Lemma 5.4.
If $L$ is a wheel, then every symmetric $2$-cycle in $L^{\times 2}_{\Delta}$ is zero.
Proof.
For subgraphs $L_{1},L_{2}\subset L$ we write $L_{2}\rightsquigarrow L_{1}$ if any edge in $L_{2}-L_{1}$ and simple cycle in $L_{2}$ have a common vertex. If $L_{2}\rightsquigarrow L_{1}$, then any symmetric $2$-cycle in $(L_{2})^{\times 2}_{\Delta}$ lies in $(L_{1})^{\times 2}_{\Delta}$. (Indeed, apply Lemma 5.3.a to the support $L^{\prime}\subset L_{2}$ of the $2$-cycle. We obtain that in $L^{\prime}$ any edge $e$ is vertex-disjoint with some simple cycle. Since $L_{2}\rightsquigarrow L_{1}$, we have $e\subset L_{1}$.)
Now the lemma follows because
(a wheel on n+1 vertices)(a cycle graph on n $(\text{a~{}wheel on $n+1$ vertices})\quad\rightsquigarrow\quad(\text{a~{}cyclegraph on $n$ vertices})\quad\rightsquigarrow\quad\varnothing.\qed$
Proof of Lemma 5.1.
If $L$ contains vertex-disjoint cycles $P,Q$, then the only non-zero symmetric $2$-cycle $C$ in $L^{\times 2}_{\Delta}$ is $P\times Q+Q\times P$. Since $L$ is the support of $C$, it follows that $L=P\sqcup Q$.
So assume that $L$ contains no two vertex-disjoint cycles.
By Lemma 5.3.b, $L$ has no vertices of degree $1$.
In this paragraph we show that it suffices to prove the lemma for the case when $L$ has no vertices of degree $2$ (then ‘homeomorphic’ is equivalent to ‘isomorphic’). Let $v$ be a vertex adjacent to exactly two vertices $v_{1},v_{2}$. By Lemma 5.3.c, there exists a simple cycle $C$ not passing through any of $v,v_{1},v_{2}$. The graph $L$ does not contain an edge $v_{1}v_{2}$: otherwise its simple cycles $C$ and $(vv_{1}v_{2})$ would be vertex-disjoint. Let $L_{0}$ be the graph obtained from $L$ by deleting the vertex $v$ and adding an edge $v_{1}v_{2}$ (then $L$ is obtained from $L_{0}$ by edge subdivision). The graph $L_{0}$ has fewer vertices than $L$, contains no two vertex-disjoint cycles, and is the support of the only non-zero symmetric $2$-cycle in $(L_{0})^{\times 2}_{\Delta}$. By continuing this procedure we obtain the required claim.
So assume that $L$ contains no two vertex-disjoint cycles, and no vertices of degrees $\leq 2$. By Dirac Theorem [Lo79, §10, ex. 4, pp. 377–378],
if a graph has no vertices of degrees $\leq 2$, and contains no two vertex-disjoint cycles, then it is isomorphic to $K_{5}$, $K_{3,3}$, or a wheel.
Hence by Lemma 5.4 $L$ is isomorphic to $K_{5}$ or $K_{3,3}$. ∎
Proof of Lemma 5.2.
Denote given $2$-cycle by $C$. Suppose to the contrary that $L^{\times 2}_{\Delta}$ contains a non-zero symmetric $2$-cycle $C^{\prime}\neq C$. Then $C^{\prime}$ contains a cell $(e_{1},e_{2})\notin C$. Construct $L_{0}$ by identifying interior points of edges $e_{1},e_{2}$ as a single vertex $a$. The edges of $L_{0}$ that are not obtained by cutting edges of $L$ are edges of $L$. Then symmetric $2$-cycles in $(L_{0})^{\times 2}_{\Delta}$ are exactly symmetric $2$-cycles in $L^{\times 2}_{\Delta}$ not containing $(e_{1},e_{2})$. Hence $C$ is an $s$-minimal $2$-cycle in $(L_{0})^{\times 2}_{\Delta}$ with support $L_{0}$. On the other hand, since $(e_{1},e_{2})\in C^{\prime}$, the set $(L_{0})^{\times 2}_{\Delta}$ contains fewer symmetric $2$-cycles than $L^{\times 2}_{\Delta}$.
If $(L_{0})^{\times 2}_{\Delta}$ contains more than one non-zero symmetric $2$-cycle, then repeat the procedure. Otherwise $L_{0}$ is as in the hypothesis of Lemma 5.1. Graphs $K_{3,3}$ and $P\sqcup Q$ for vertex-disjoint cycles $P,Q$ have no vertices of degree $4$. Since $\deg_{L_{0}}a=4$, by Lemma 5.1 it follows that $L_{0}$ is homeomorphic to $K_{5}$. Hence $L$ is homeomorphic to the graph $L^{\prime}$ obtained from the graph $K_{4}$ on vertices $v_{i},v^{\prime}_{i}$ ($i=1,2$) by adding new vertices $a_{i}$ ($i=1,2$) and edges $a_{i}v_{i},a_{i}v^{\prime}_{i}$ ($i=1,2$). Since the graph $L^{\prime}-v_{1}-v_{2}$ is a simple path $a_{1}v^{\prime}_{1}v^{\prime}_{2}a_{2}$, the graph $L^{\prime}$ does not contain a cycle vertex-disjoint with the edge $v_{1}v_{2}$. This contradicts Lemma 5.3.a. ∎
6 Appendix: some alternative proofs
A direct proof of Lemma 2.1.b for $k=1$.
Denote by $\alpha(a,b;u,v,w)$ and $\beta(a,b,c;u,v)$ the relations from the statement. For any $b,u,v\in[n]-\{1\}$ such that $u\neq v$, one has
for any $\text{for any }a\in[n]-\{b\}$ $au^{\prime}bv^{\prime}=a1^{\prime}bu^{\prime}+a1^{\prime}bv^{\prime}+\alpha(a,b;1,u,v);$
for any $\text{for any }a\in[n]-\{1,b\}$ $a1^{\prime}bv^{\prime}=11^{\prime}av^{\prime}+11^{\prime}bv^{\prime}+\beta(1,a,b;1,v).$
Hence any linear relation on cycles of length 4 in $K_{n,n}$ can be reduced (by adding some of $\alpha$’s and $\beta$’s) to a relation including only cycles $11^{\prime}bv^{\prime}$. The latter relation is trivial because for any $a,b,u,v\in[n]-\{1\}$ the edge $\{a,u^{\prime}\}$ belongs to the cycle $11^{\prime}bv^{\prime}$ if and only if $a=b$ and $u=v$. ∎
Alternative proof of Theorem 2.3.b.
Induction on $\ell$.
Let us prove the base $\ell=1$. If $n=3$, then $H_{1}(\widetilde{K_{3}})=\mathop{Ker}(I+t)=\{0,\widetilde{K_{3}}\}$ and $\mathop{Im}(I+t)=0$. Assume further that $n\geq 4$. Observe that $\widetilde{K_{n}}/t=K_{n}$. Denote by $q:\widetilde{K_{n}}\to K_{n}$ the quotient map. Since $t$ has no fixed points, the restriction of the induced map $q_{*}:H_{1}(\widetilde{K_{n}})\to H_{1}(K_{n})$ to the subgroup $\mathop{Ker}(I+t)\subset H_{1}(\widetilde{K_{n}})$ of $t$-symmetric cycles is an isomorphism. (Indeed, the inverse map $R:H_{1}(K_{n})\to\mathop{Ker}(I+t)$ is defined by $R(y):=q^{-1}(y)$.) Clearly, $q(\widetilde{K_{3}})=(1,2,3)$, and for a cycle $m$ of length 4 in $\widetilde{K_{n}}$ we have $q(m+tm)=q(m)$ is a cycle of length 4 in $K_{n}$. Hence by Theorem 2.3.a and Lemma 2.1.a $q\mathop{Im}(I+t)$ is generated by cycles of length 4 in $K_{n}$.
In the following paragraph we prove that any $1$-cycle in $K_{n}$ is a sum of some cycles of length $4$ and, possibly, of the cycle $(1,2,3)$.
The sum of two cycles of length $3$ having a common edge is a cycle of length $4$. So if a cycle $\alpha$ of length $3$ is a sum of $(1,2,3)$ and cycles of length 4, then any cycle of length $3$ having a common edge with $\alpha$ is such. Hence all cycles of length $3$ are such. Now the required assertion follows because by Statement 2.4 any $1$-cycle in $K_{n}$ is a sum of some cycles of length 3.
Then $\mathop{Ker}(I+t)$ is generated by $\widetilde{K_{3}}$ and $\mathop{Im}(I+t)$. The number of edges modulo 4 defines a homomorphism $\mathop{Ker}(I+t)\to{\mathbb{Z}}_{4}$. This homomorphism is 2 on $\widetilde{K_{3}}$ and is 0 on $\mathop{Im}(I+t)$. Hence $\widetilde{K_{3}}\not\in\mathop{Im}(I+t)\subset H_{1}(\widetilde{K_{n}})$.
The inductive step $\ell\to\ell+1$ is obtained by applying the following Lemma 6.1 to
$W_{1}=H_{1}(\widetilde{K_{n}}),\quad W_{2}=H_{1}(\widetilde{K_{n}})^{\otimes\ell},\quad x_{1}=\widetilde{K_{3}},\quad x_{2}=\widetilde{K_{3}}^{\otimes\ell},\quad A_{1}=t,\quad A_{2}=t^{\otimes\ell}.$
Lemma 6.1.
For $j=1,2$ let $W_{j}$ be a vector space over ${\mathbb{Z}}_{2}$,  $x_{j}\in W_{j}-\{0\}$ a vector, and $A_{j}$ an operator in $W_{j}$ such that $\mathop{Ker}(I+A_{j})=\mathop{Im}(I+A_{j})\oplus\left\langle x_{j}\right\rangle$. Then the operator $A:=A_{1}\otimes A_{2}$ in $W:=W_{1}\otimes W_{2}$ satisfies $\mathop{Ker}(I+A)=\mathop{Im}(I+A)\oplus\left\langle x_{1}\otimes x_{2}\right\rangle$.
Proof.
Let $N_{j}:=I+A_{j}$. Then $\mathop{Ker}N_{j}=(\mathop{Im}N_{j})\oplus\left\langle x_{j}\right\rangle$. Hence $N_{j}^{2}=0$.
Denote by $J_{m}(\lambda)$ the Jordan cell of size $m$ with eigenvalue $\lambda$. Take the Jordan form of $N_{j}$, i.e., represent $W_{j}$ as a direct sum of $N_{j}$-invariant subspaces $W_{j,p}$, $0\leq p\leq m_{j}$, such that
$\bullet$ $W_{j,0}=\left\langle x_{j}\right\rangle$, $N_{j}W_{j,0}=0$;
$\bullet$ for every $p>0$, the subspace $W_{j,p}$ is two-dimensional, and the operator $N_{j}$ acts on it with matrix $J_{2}(0)$ (in some basis).
Hence the operator $A_{j}$ also preserves every subspace $W_{j,p}$, acting on it identically if $p=0$, and with matrix $J_{2}(1)$ (in some basis) otherwise. So the space $W$ is a direct sum of $A$-invariant subspaces $W_{p_{1},p_{2}}:=W_{1,p_{1}}\otimes W_{2,p_{2}}$, $0\leq p_{j}\leq m_{j}$. Now the lemma holds because $W_{0,0}=\left\langle x_{1}\otimes x_{2}\right\rangle$, $(I+A)W_{0,0}=0$, and
(*)   $\mathop{Ker}(I+A|_{U})=\mathop{Im}(I+A|_{U})$ on every other summand $U:=W_{p_{1},p_{2}}$.
Let us prove assertion (*). Denote by $M(X)$ the matrix of an operator $X$ in the considered basis.
If either $p_{1}>0=p_{2}$ or $p_{2}>0=p_{1}$, then $U$ has a basis $e_{1},e_{2}$ in which $M(A|_{U})=J_{2}(1)$. So $M(I+A|_{U})=J_{2}(0)$. Thus, $\mathop{Ker}(I+A|_{U})=\mathop{Im}(I+A|_{U})=\left\langle e_{1}\right\rangle$.
If $p_{1},p_{2}>0$, then $U$ has a basis $e_{1},e_{2},e_{3},e_{4}$ in which
$M(A|_{U})=\begin{pmatrix}J_{2}(1)&J_{2}(1)\\ 0&J_{2}(1)\end{pmatrix}=\begin{pmatrix}1&1&1&1\\ 0&1&0&1\\ 0&0&1&1\\ 0&0&0&1\end{pmatrix},\quad\text{so}\quad M(I+A|_{U})=\begin{pmatrix}0&1&1&1\\ 0&0&0&1\\ 0&0&0&1\\ 0&0&0&0\end{pmatrix}.$
Hence $\mathop{Ker}(I+A|_{U})=\mathop{Im}(I+A|_{U})=\left\langle e_{1},e_{2}+e_{3}\right\rangle$. ∎
Remark 6.2.
Denote by $A^{\prime}$ a copy of $A$. The simplicial deleted join $G^{*2}_{\Delta}$ of a complex $G=(V,F)$ is the complex with the set $V\sqcup V^{\prime}$ of vertices and the set $\{\sigma*\tau^{\prime}\ :\ \sigma,\tau\in F,\ \sigma\cap\tau=\varnothing\}$ of faces (here $\sigma$ and $\tau$ are considered as faces of a simplex, so that $\sigma*\tau^{\prime}$ is the face spanned by all the vertices of $\sigma$ and of $\tau^{\prime}$; if $\sigma$ and $\tau$ are considered as subsets, then $\sigma*\tau^{\prime}$ should be replaced by $\sigma\sqcup\tau^{\prime}$).
Analogously to the isomorphism (1) from the proof of Theorem 3.1 one proves that for complexes $A,B$ of dimensions $a,b$ the map $(A*B)^{\times 2}_{\Delta}\to A^{*2}_{\Delta}\times B^{*2}_{\Delta}$ given by $(\alpha_{1}*\beta_{1})\times(\alpha_{2}*\beta_{2})\mapsto(\alpha_{1}*\alpha_{2})\times(\beta_{1}*\beta_{2})$ on $(a+b+2)$-cells induces an isomorphism of the groups $H_{a+b+2}$ of cellular $(a+b+2)$-cycles. This is the analogue of [Ma03, Lemma 5.5.2] for deleted products.
An alternative definition of the isomorphism (1) is the composition
$H_{2}((K_{n,n})^{\times 2}_{\Delta})\overset{(a)}{\cong}H_{3}((K_{n,n})^{*2}_{\Delta})\overset{(b)}{\cong}H_{3}(\widetilde{K_{n}}*\widetilde{K_{n}})\overset{(c)}{\cong}H_{2}(\widetilde{K_{n}}\times\widetilde{K_{n}}),\quad\text{where}$
$\bullet$ isomorphism (a) holds by [PS20, Lemma 4d], cf. [Sk, 7.2.8c and 7.2.3b]; it is induced by the map $\sigma\times\tau\mapsto\sigma*\tau$ from the set of $2$-cells to the set of $3$-faces;
$\bullet$ isomorphism (b) holds because $K_{n,n}=[n]^{*2}$ and $\widetilde{K_{n}}=[n]^{*2}_{\Delta}$, so $(K_{n,n})^{*2}_{\Delta}\cong\widetilde{K_{n}}*\widetilde{K_{n}}$ [Ma03, Lemma 5.5.2]; the latter homeomorphism maps $(\alpha_{1}*\alpha_{2})*(\beta_{1}*\beta_{2})$ to $(\alpha_{1}*\beta_{1})*(\alpha_{2}*\beta_{2})$ for $\alpha,\beta\in[n]^{2}$ such that $\alpha_{j}\neq\beta_{j}$ for every $j\in[2]$;
$\bullet$ isomorphism (c) is induced by the map $\sigma*\tau\mapsto\sigma\times\tau$ from the set of $3$-faces to the set of $2$-cells.