### v4

# Embeddability of joinpowers

Abstract
A general position map $f:K\to M$ of a $k$-dimensional simplicial complex to a $2k$-dimensional manifold (for $k=1$, of a graph to a surface) is a ${\mathbb{Z}}_{2}$-embedding if $|f\sigma\cap f\tau|$ is even for any vertex-disjoint $k$-dimensional faces (for $k=1$, edges) $\sigma,\tau$. We present criteria for ${\mathbb{Z}}_{2}$-embeddability of certain $k$-dimensional complex (for $k=1$, of any graph) to a $2k$-dimensional manifold. These criteria are
$\bullet$ a ‘Kuratowski-type’ version of the Fulek-Kynčl-Bikeev criteria (for $k=1$), and
$\bullet$ a converse to the Dzhenzher-Skopenkov necessary condition (for $k>1$).
Our higher-dimensional criterion allows us to reduce the modulo 2 Kühnel problem on embeddings to a purely algebraic problem.
Our proof is interplay between geometric topology, combinatorics and linear algebra. It is based on calculation of generators in the homology of certain configuration space (the deleted product) of certain complex (joinpower).
MSC 2020: 57Q35, 55N91, 05E45.
Keywords: embedding, mod 2 embedding, joinpower, deleted product, homology.
1 Introduction and main results
Description of main results
We abbreviate ‘$k$-dimensional’ to just ‘$k$-’.
In this paper a complex is the set of some closed faces of some simplex (a longer term is finite simplicial complex). We identify this set with the union of those faces (i.e. with the body of the complex). A $k$-complex is the union of some closed faces of dimension at most $k$ (for $k=1$ this is a graph).
A general position map $K\to M$ of a $k$-complex to a $2k$-manifold (see the definition e.g. in [Sk24, §1.1]) is called a ${\mathbb{Z}}_{2}$-embedding if $|f\sigma\cap f\tau|$ is even for any vertex-disjoint $k$-faces $\sigma,\tau$.
We present criteria for ${\mathbb{Z}}_{2}$-embeddability of certain $k$-complex (for $k=1$, of any graph) to $2k$-manifolds.
These criteria are
$\bullet$ a ‘Kuratowski-type’ version of the Fulek-Kynčl-Bikeev criteria, i.e., a version involving only obstructions coming from $K_{5}$ and $K_{3,3}$ (for $k=1$; Theorems 1.1, 1.2 and 1.3), and
$\bullet$ a converse to the Dzhenzher-Skopenkov necessary condition (for $k>1$; Theorem 1.5).
Our higher-dimensional criterion allows us to reduce the modulo 2 Kühnel problem on embeddings to a purely algebraic problem (Conjecture 1.7).
Our proof is an interplay between geometric topology, combinatorics, and linear algebra. It is based on the calculation of generators in the homology of a certain configuration space (the deleted product) of a certain complex (joinpower), see Theorems 3.1, 3.2, and Remark 1.6.
Some motivation
This subsection is formally not used below.
The classical Heawood inequality provides a lower estimation for a number $g$ such that the complete graph $K_{n}$ on $n$ vertices embeds into the sphere with $g$ handles. A higher-dimensional analogue of the Heawood inequality is the Kühnel conjecture on embeddings. In a simplified form it asks for an analogous estimation of the ‘complexity’ (i.e., of $\dim H_{k}(M;{\mathbb{Z}}_{2})$) of a $2k$-manifold $M$ into which the union of all $k$-faces of $n$-simplex is embedded. See the detailed discussion of the conjecture in [DS22, Remark 1.3], cf. [Ku23, §4].
Recent results on the conjecture work not only for embeddings but for ${\mathbb{Z}}_{2}$-embeddings ([PT19, Theorem 1], [DS22]; see [KS21] for a simplified direct proof of [PT19, Theorem 1]). These results involve a necessary condition for ${\mathbb{Z}}_{2}$-embeddability of certain $k$-complexes to $2k$-manifolds (the ‘only if’ parts of Theorems 1.1, 1.2 and 1.5; similarly one proves the ‘only if’ part of Theorem 1.3, which part was not stated). This condition is stated (implicitly in [PT19] and explicitly in [DS22, KS21]) in terms of finding the minimal rank of a matrix with certain relations on its entries, and having a certain special form. So this condition is related to the low rank matrix completion problem, and to the Netflix problem from machine learning. Our main result is the sufficiency of this condition (the ‘if’ parts of Theorems 1.11.3 and 1.5). See survey [DGN+] and the references therein.
Known results on embeddability of $k$-complexes in $2k$-manifolds are described in [Sk24, Remark 1.1.7.bc]. Some motivations for studies of ${\mathbb{Z}}_{2}$-embeddability are presented in [Sk24, §1.2]. The particular case of ${\mathbb{Z}}_{2}$-embeddability of graphs to surfaces is actively studied, see survey [Sc13] and recent papers [Ky16, FK19, Bi21].
Our main instrument is a description of generators in the homology of certain configuration space (deleted product), see Theorem 3.2 (and Theorem 3.1 due to [Ho07]). For related results
$\bullet$ on graphs see [Ni00, ST03, Ho07, BF09, Ho09, FH10], [KP11, Theorem 3.16], [Ky16, HNT, AK20, Dz25], and references in [AK20];
$\bullet$ on higher-dimensional complexes see [Di08, §19], [Me06, Theorem 3.8].
(References to papers published earlier than 2000 can be found in the above papers.)
Notation and conventions
Below all matrices have ${\mathbb{Z}}_{2}$-entries. Assume that $n\geq 3$. Let $n]:=\{1,2,\ldots,n\}$.
Denote by $K_{n,n}$ the complete bipartite graph with parts $[n]$ and a copy $n]^{\prime}:=\{k^{\prime}\ :\ k\in[n]\}$ of $[n]$.
Denote by $I_{n}$ the identity matrix of size $n$. Denote by $H_{n}$ the square matrix of size $2n$ whose off-diagonal $2\times 2$-blocks are zero matrices, and whose diagonal $2\times 2$-blocks are $\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$. For a matrix $A$ let $A_{P,Q}$ be the entry in the row of $P$ and column of $Q$.
A (simplicial) $1$-cycle (modulo 2) in a graph $K$ is a set $C$ of edges such that every vertex is contained in an even number of edges from $C$. A (simplicial) $k$-cycle (modulo 2) in a complex $K$ is a set $C$ of $k$-faces such that every (k1)(k-1)-face is contained in an even number of faces from $C$.
We omit ${\mathbb{Z}}_{2}$-coefficients of homology groups considered below. Denote by $H_{k}(K)$ the group of all $k$-cycles in a $k$-complex $K$, with the sum modulo 2 operation. For a $2k$-manifold $M$ denote by $H_{k}(M)$ the modulo 2 homology group, and by $\cap_{M}:H_{k}(M)\times H_{k}(M)\to{\mathbb{Z}}_{2}$ the modulo 2 intersection form (for definitions accessible to non-specialists see [Sk20, §6, §10]).
Results for graphs
Our results for graphs are Theorems 1.1, 1.2 and 1.3 (of which Theorem 1.1 is a particular case of Theorem 1.3, and Theorem 1.2 is a particular case of both Theorems 1.3 and 1.5).
Let $M_{m}$ be the sphere with $m$ Möbius films, $S_{g}$ the sphere with $g$ handles, and $N$ either $M_{m}$ or $S_{g}$. Denote
$\Omega_{M_{m}}:=I_{m},\quad\beta(M_{m}):=m,\quad\Omega_{S_{g}}:=H_{g},\quad\text{and}\quad\beta(S_{g}):=2g.$
Let $A$ be a symmetric square matrix of $\binom{n}{3}$ rows (whose rows and whose columns correspond to all 3-element subsets of $[n]$). The matrix $A$ is said to be
$\bullet$ independent if $A_{P,Q}=0$ whenever $P\cap Q=\varnothing$;
$\bullet$ additive if
$A_{F-a,Q}+A_{F-b,Q}+A_{F-c,Q}+A_{F-d,Q}=0$
for any 4-element subset $F=\{a,b,c,d\}\subset[n]$ and 3-element subset $Q\subset[n]$ (here we abbreviate $\{a\}$ to $a$).
$\bullet$ non-trivial if for any $5$-element subset $F\subset[n]$ and [^3] Here and below after Theorem 1.1
$\bullet$ ‘for any vertex $v$ / edge $e$’ is equivalent to ‘for some vertex $v$ / edge ee[PT19, Lemma 20], [KS21, Proposition 2.2 and Remark 2.3], [DS22, Remark 4.4];
$\bullet$ strictly speaking, the properties of matrix $A$ should be called $K_{5}$-independence’, $K_{3,3}$-independence’, etc (for the sake of readability, we shorten the notation).
$v\in F$ we have $S_{F,v}A=1$, where $S_{F,v}A$ is the sum of $A_{P,Q}$ over all the $3$ unordered pairs $\{P,Q\}$ of 33-element subsets $P,Q\subset F$ such that $P\cap Q=\{v\}$.
Theorem 1.1.
A ${\mathbb{Z}}_{2}$-embedding $K_{n}\to N$ exists if and only if there is a $\beta(N)\times{n\choose 3}$ matrix $Y$ such that $Y^{T}\Omega_{N}Y$ is an independent additive non-trivial matrix.
Let $A$ be a symmetric square matrix of $\binom{n}{2}^{2}$ rows (whose rows and whose columns correspond to all cycles of length 4 in $K_{n,n}$). The matrix $A$ is said to be
$\bullet$ independent if $A_{P,Q}=0$ whenever $P$ and $Q$ are vertex-disjoint;
$\bullet$ additive if $A_{P,Q}=A_{X,Q}+A_{Y,Q}$ whenever $P=X\oplus Y$ (as sets $P,X,Y$ of edges).
$\bullet$ non-trivial if for any subgraph $X$ of $K_{n,n}$ isomorphic to $K_{3,3}$ and edge $e$ of $X$ we have $S_{X,e}A=1$, where $S_{X,e}A$ is the sum of $A_{P,Q}$ over all the 2 unordered pairs $\{P,Q\}$ of cycles $P,Q\subset X$ of length 4 whose only common edge is $e$.
Theorem 1.2.
A ${\mathbb{Z}}_{2}$-embedding $K_{n,n}\to N$ exists if and only if there is a $\beta(N)\times{n\choose 2}^{2}$ matrix $Y$ such that $Y^{T}\Omega_{N}Y$ is an independent additive non-trivial matrix.
Theorem 1.3.
A ${\mathbb{Z}}_{2}$-embedding $K\to N$ of a graph $K$ exists if and only if there is a homomorphism $y:H_{1}(K)\to H_{1}(N)$ such that
(independence) $y(P)\cap_{N}y(Q)=0$ whenever $P$ and $Q$ are vertex-disjoint;
($K_{5}$-non-triviality) for any subgraph $F$ homeomorphic to $K_{5}$, and vertex $v$ of $F$ we have $S_{F,v}y=1$, where $S_{F,v}y$ is the sum of $y(P)\cap_{N}y(Q)$ over all the 3 unordered pairs $\{P,Q\}$ of cycles $P,Q\subset F$ of length [^4] Here and in ($K_{3,3}$-non-triviality) the cycles of length 3 and 4 are not the cycles of that length in $K$, but rather they are cycles with 3 and 4, respectively, vertices of degree more than 2 in $F$ and in $X$. 3 such that $v$ is the only common vertex of $P$ and $Q$; and
($K_{3,3}$-non-triviality) for any subgraph $X$ homeomorphic to $K_{3,3}$, and edge $e$ of $X$ we have $S_{X,e}y=1$, where $S_{X,e}y$ is the sum of $y(P)\cap_{N}y(Q)$ over all the 2 unordered pairs $\{P,Q\}$ of cycles $P,Q\subset X$ of length 4 such that $e$ is the only common edge of $P$ and $Q$.
Remark 1.4.
(a) (on the novelty and idea of proof) The ‘only if’ part of Theorem 1.3 is essentially known, and is analogous to [DS22, Remark 1.6], which in turn is analogous to [FK19, Bi21]. Proof of the ‘if’ part is analogous to the proof of the ‘if’ part of Theorem 1.5 for $k=1$ in §4, except that reference to Theorem 3.2.b is replaced by reference to Theorem 3.1. See more comments on the proof in Remark 1.6.c.
(b) In matrix form, the condition of Theorem 1.3 means that there are a set of $r$ generators of $H_{1}(K)$, and a $\beta(N)\times r$ matrix $Y$ such that $A:=Y^{T}\Omega_{N}Y$ is independent additive non-trivial (as defined below).
Let $A$ be a symmetric $r\times r$ matrix whose rows and whose columns correspond to some generators of $H_{1}(K)$. The matrix $A$ is said to be
$\bullet$ independent, if $A_{P,Q}=0$ whenever $P$ and $Q$ are vertex-disjoint;
$\bullet$ additive, if $A_{P,Q}=A_{P_{1},Q}+\ldots+A_{P_{t},Q}$ whenever $P=P_{1}\oplus\ldots\oplus P_{t}$ (as sets of edges).
If $X,Y\in H_{1}(K)$, then express them as a sum of the generators, $X=X_{1}+\ldots+X_{p}$, $Y=Y_{1}+\ldots+Y_{q}$, and define $A_{X,Y}:=\sum_{i,j}A_{X_{i},Y_{j}}$. The matrix $A$ is said to be $K_{5}$-non-trivial ($K_{3,3}$-non-trivial), if the $K_{5}$-non-triviality ($K_{3,3}$-non-triviality) of Theorem 1.3 holds replacing $y(P)\cap_{N}y(Q)$ by $A_{P,Q}$.
The main result for higher dimensions
We consider piecewise-linear (PL) manifolds as defined in [RS72]. Let $M$ be a closed $2k$-manifold. Denote $\beta(M):=\dim H_{k}(M)$. Denote $\Omega_{M}:=I_{\beta(M)}$ if there is $x\in H_{k}(M)$ such that $x\cap_{M}x=1$, and $\Omega_{M}:=H_{\beta(M)/2}$ otherwise, when $\beta(M)$ is known to be even ($\Omega_{M}$ is the matrix of $\cap_{M}$ in some basis).
Denote $D^{j}:=[0,1]^{j}$ and $S^{j-1}:=\partial D^{j}$. A complex $M$ is called $m$-connected if for any $j=0,1,\ldots,m$ every continuous map $f:S^{j}\to M$ extends continuously over $D^{j+1}$.
Let $n]^{*k+1}$ be the $k$-complex
$\bullet$ whose vertex set is the $(k+1)\times n$ grid $\{0,\ldots,k\}\times[n]$,
$\bullet$ in which every $k+1$ vertices from different rows (i.e. with different first coordinates) span a $k$-simplex.
For $k=1$ this is $K_{n,n}$. For geometric interpretation see [Ma03, Proposition 4.2.4]. A $k$-octahedron is the set of $k$-faces of a subcomplex (of $n]^{*k+1}$) isomorphic to $2]^{*k+1}\cong S^{k}$. For $2$-element subsets $P_{0},\ldots,P_{k}\subset[n]$ such a subcomplex is defined by the set $0\times P_{0}\sqcup\ldots\sqcup k\times P_{k}$ of its vertices.
Let $A$ be a symmetric square matrix of $\binom{n}{2}^{k+1}$ rows (whose rows and whose columns correspond to all $k$-octahedra). The matrix $A$ is said to be
$\bullet$ independent if $A_{P,Q}=0$ whenever $P$ and $Q$ are vertex-disjoint;
$\bullet$ additive if $A_{P,Q}=A_{X,Q}+A_{Y,Q}$ whenever $P=X\oplus Y$ (as sets $P,X,Y$ of $k$-faces);
$\bullet$ non-trivial if for any subcomplex $X$ of $n]^{*k+1}$ isomorphic to $3]^{*k+1}$ and $k$-face $e$ of $X$ we have $S_{X,e}A=1$, where by $S_{X,e}A$ we denote the sum of $A_{P,Q}$ over all $2^{k}$ unordered pairs $\{P,Q\}$ of $k$-octahedra $P,Q\subset X$ whose only common $k$-face is $e$.
Theorem 1.5.
Let $M$ be a closed (k1)(k-1)-connected $2k$-manifold. A ${\mathbb{Z}}_{2}$-embedding $n]^{*k+1}\to M$ exists if and only if there is a $\beta(M)\times{n\choose 2}^{k+1}$ matrix $Y$ such that $Y^{T}\Omega_{M}Y$ is an independent additive non-trivial matrix.
Remark 1.6.
(a) (on the novelty and idea of proof) The ‘only if’ part of Theorem 1.5 is [DS22, Remark 1.9b]. Our main achievement is the ‘if’ part (§4). For this we use a homological ${\mathbb{Z}}_{2}$-embeddability criterion (Theorem 4.3 by E. Kogan; for completeness, we present in §4 a standard deduction from known results). Our main idea is that the independence and non-triviality of the matrix $Y^{T}\Omega_{M}Y$ correspond to generators of the homology group of a certain configuration space (Theorem 3.2).
(b) A version of Theorem 1.5 for the $k$-skeleton of $n$-simplex instead of $n]^{*k+1}$ is stated analogously (cf. Theorem 1.1). Such a version is presumably false for $k>1$ because the relevant homology group of configuration space is complicated. So Theorem 1.5 is not as trivial a corollary of Theorem 4.3 as it may seem.
(c) Theorem 1.5 generalizes to joins of finite sets of different sizes $n_{0},\ldots,n_{k}\geq 3$: a ${\mathbb{Z}}_{2}$-embedding $n_{0}]*\ldots*[n_{k}]\to M$ exists if and only if there is a $\beta(M)\times{n_{0}\choose 2}\ldots{n_{k}\choose 2}$ matrix $Y$ such that $Y^{T}\Omega_{M}Y$ is an independent additive non-trivial matrix. The proof of Theorem 1.5 generalizes trivially to a proof of the generalization, see Remarks 2.6, 3.5, 4.6. In those Remarks, the changes of the proofs that repeat changes of the statements are omitted.
Conjecture 1.7.
(a) (asymptotic mod 2 Kühnel conjecture) For any integer $k>0$ there is $c_{k}>0$ such that if $n]^{*k+1}$ has a ${\mathbb{Z}}_{2}$-embedding to a closed $2k$-manifold $M$, then $\beta(M)\geq c_{k}n^{k+1}$.
By Theorem 1.5, this is equivalent to the following algebraic conjecture:
For any integer $k>0$ there is $c_{k}>0$ such that if there is a $d\times{n\choose 2}^{k+1}$ matrix $Y$ such that either $Y^{T}Y$ or $Y^{T}H_{d/2}Y$ is independent additive non-trivial, then $d\geq c_{k}n^{k+1}$.
Both conjectures are proved for $k=1$ [FK19], and for any $k$ replacing $n^{k+1}$ by $n^{2}$ [DS22].
(b) (analogue of Theorem 1.5 for embeddings) Assume that $k\geq 3$ and $M$ is a closed (k1)(k-1)-connected $2k$-manifold. An embedding $n]^{*k+1}\to M$ exists if and only if there is a $\beta(M)\times{n\choose 2}^{k+1}$ matrix $Y$ with integer entries such that $Y^{T}\Omega_{M,{\mathbb{Z}}}Y$ is an independent additive non-trivial matrix.
Here $\Omega_{M,{\mathbb{Z}}}$ is the matrix of the intersection form of $M$ in some basis of $H_{k}(M;{\mathbb{Z}})\cong{\mathbb{Z}}^{\beta(M)}$ (the isomorphism holds since $M$ is (k1)(k-1)-connected).
Let $A$ be a $(-1)^{k}$-symmetric square matrix with integer entries, of $\binom{n}{2}^{k+1}$ rows. Independence is defined analogously to the mod 2 case. The matrix $A$ is said to be
$\bullet$ additive if $A_{P,Q}=A_{X,Q}+A_{Y,Q}$ whenever $P=X+Y$ (as integer $k$-cycles);
$\bullet$ non-trivial if for any subcomplex $X$ of $n]^{*k+1}$ isomorphic to $3]^{*k+1}$ and $k$-face $e$ of $X$ the sum $S_{X,e}A$ is odd.
The ‘only if’ part of this conjecture is proved analogously to [DS22, Theorem 1.8 and Remark 1.9b].
2 Generators in homology of the complex
Of this section, only Lemmas 2.1, 2.3 and 2.5 are used later.
Lemma 2.1.
(a) Any $k$-cycle in $n]^{*k+1}$ is a sum of some $k$-octahedra.
(b) Any linear relation on $k$-octahedra is a sum of some relations
$K*\{u,v\}+K*\{v,w\}+K*\{w,u\}=0$
for pairwise distinct $u,v,w\in[n]$, and (k1)(k-1)-octahedra $K$, and relations obtained from that by permutations of ‘join coordinates’ [^5] By * we denote a join. Recall that a (k1)(k-1)-octahedron is the join $K=K_{0}*\ldots*K_{k-1}$ of some 2-element subsets $K_{0},\ldots,K_{k-1}\subset[n]$. The ‘join coordinates’ of $K*\{a,b\}$ are $K_{0},\ldots,K_{k-1},\{a,b\}$. We take the same permutation of ‘join coordinates’ of $K*\{u,v\}$, of $K*\{v,w\}$, and of $K*\{w,u\}$. [^6] Rigorously, a linear relation is a set of $k$-octahedra such that every $k$-face of $n]^{*k+1}$ is contained in an even number of $k$-octahedra from this set..
Comment. For $k=1$ this lemma states that
(a) any $1$-cycle in $K_{n,n}$ is a sum of some cycles of length $4$;
(b) any linear relation on such cycles is a sum of some relations $au^{\prime}bv^{\prime}+au^{\prime}bw^{\prime}+av^{\prime}bw^{\prime}=0$ and symmetric to them relations $au^{\prime}bv^{\prime}+au^{\prime}cv^{\prime}+bu^{\prime}cv^{\prime}=0$.
Of these, (a) is clear (by Statement 2.4) and is known.
Lemma 2.1 holds by the following ‘dual’ restatement, Lemma 2.2. We use the following analogue of the duality between the faces of the regular octahedron and the vertices of the cube.
All $k$-faces of the complex $n]^{*k+1}$ are in 1–1 correspondence (duality) with vertices of the $(k+1)$-cube $n]^{k+1}$. A parallelepiped is a subset $P_{0}\times\ldots\times P_{k}\subset[n]^{k+1}$, where $P_{i}$ are 2-element subsets of $[n]$. This is dual to a $k$-octahedron.
coordinate line is a subset of $n]^{k+1}$ given by fixing all coordinates except one. Two $k$-faces of $n]^{*k+1}$ have a common face if and only if the two corresponding vertices of $n]^{k+1}$ belong to one coordinate line. A rook cycle is a subset of $n]^{k+1}$ containing an even number of vertices from every coordinate line. This is dual to a $k$-cycle in $n]^{*k+1}$. The property of being a rook cycle is preserved under modulo $2$ summation of subsets.
Lemma 2.2.
(a) Any rook cycle is a sum of some parallelepipeds.
(b) Any linear relation on parallelepipeds in $n]^{k+1}$ is a sum of some relations
$P\times\{a,b\}+P\times\{b,c\}+P\times\{c,a\}=0$
for pairwise distinct $a,b,c\in[n]$, and parallelepipeds $P\subset[n]^{k}$, and relations obtained from that by permutations of coordinates [^7] Rigorously, a linear relation is a set of parallelepipeds such that every vertex of $n]^{k+1}$ is contained in an even number of parallelepipeds from this set..
Proof.
For $a\in[n-1]^{k+1}$ denote by $P(a):=\{n,a_{0}\}\times\ldots\times\{n,a_{k}\}$ the parallelepiped with opposite vertices $a$ and $(n,\ldots,n)$.
For (a) it suffices to prove that any rook cycle $C\subset[n]^{k+1}$ equals to the sum $P(C)$ of parallelepipeds $P(a)$ over $a\in C\cap[n-1]^{k+1}$. Let us prove this. The sum $C+P(C)$ is a rook cycle. Since $P(a)\cap[n-1]^{k+1}=\{a\}$, we have $(C+P(C))\cap[n-1]^{k+1}=\varnothing$. Together with any vertex $c$ whose $j$-th coordinate is $n$, the sum $C+P(C)$ contains a vertex $c^{\prime}$ having the same coordinates except the $j$-th coordinate, and having fewer coordinates equal to $n$ than $c$. Hence if non-empty, $C+P(C)$ contains an element from $n-1]^{k+1}$. Thus $C+P(C)$ is empty, i.e., $C=P(C)$.
Part (b) is the following statement for $r=k+1$. For any parallelepiped $P=P_{0}\times\ldots\times P_{k}$ define the complexity to be the number of indices $j$ such that $P_{j}\subset[n-1]$. Recall footnote 7.
Every linear relation on parallelepipeds of complexities at most $r$ is a sum of some relations from the conclusion of (b).
Let us prove this statement by induction on $r$.
If $r=0$, then there are pairwise distinct vertices $a^{1},\ldots,a^{s}\in[n-1]^{k+1}$ such that the given relation is $P(a^{1})+\ldots+P(a^{s})=0$. The relation has no summands because
$\varnothing=(P(a^{1})+\ldots+P(a^{s}))\cap[n-1]^{k+1}=\{a^{1},\ldots,a^{s}\}.$
Here the second equality holds because $P(a)\cap[n-1]^{k+1}=\{a\}$.
Let us prove the inductive step $r\to r+1$.
Let $P=P_{0}\times\ldots\times P_{k}$ be a parallelepiped such that $P_{j}\subset[n-1]$ for some $j$. Take $a,b\in[n-1]$ such that $P_{j}=\{a,b\}$. Let $P^{a}$ and $P^{b}$ be the parallelepipeds obtained from $P$ by replacing $P_{j}$ with $\{a,n\}$ and with $\{b,n\}$, respectively. Then one of the relations from the conclusion of (b) is $\{P,P^{a},P^{b}\}$ (written in the form $P+P^{a}+P^{b}=0$). The complexities of $P^{a}$ and $P^{b}$ are smaller than the complexity of $P$.
Consider any linear relation on parallelepipeds of complexities at most $r+1$. As in the previous paragraph, every parallelepiped of complexity $r+1$ can be replaced with parallelepipeds of complexities at most $r$, by adding a relation from the conclusion of (b). Then by the inductive hypothesis the linear relation is a sum of some relations from the conclusion of (b). ∎
Let $n]^{*2}_{\Delta}$ be the graph obtained from $K_{n,n}$ by deleting all ‘diagonal’ edges $jj^{\prime}$, $j\in[n]$. E.g. $3]^{*2}_{\Delta}$ is the cycle $12^{\prime}31^{\prime}23^{\prime}$ of length $6$. (See the sense of the subscript in a notation of a deleted join in [Ma03, §5].)
Lemma 2.3.
If $n\geq 4$, then any $1$-cycle in $n]^{*2}_{\Delta}$ is a sum of some cycles of length $4$.
For a proof we need the following simple (and so known) statement.
Statement 2.4 (simple and known).
Any $1$-cycle in a graph is a sum of some simple cycles for which there are no chords, i.e., edges in the graph between two non-consecutive vertices in the cycle.
Proof.
Any 1-cycle is a sum of some simple cycles. Any chord for a simple cycle of length $l$ gives a representation of the cycle as the sum of two simple cycles of lengths less than $l$. ∎
Proof of Lemma 2.3.
A simple cycle in $n]^{*2}_{\Delta}$ is formed by an $l$-element cyclic sequence of vertices such that
$\bullet$ the number $l$ is even and greater than $2$;
$\bullet$ the parts of vertices alternate in the sequence;
$\bullet$ within each separate part, the vertices are not repeated;
$\bullet$ there are no two consecutive vertices $m$ and $m^{\prime}$.
By Statement 2.4 we may assume that given 1-cycle is a simple cycle for which there are no chords. Then any two non-consecutive vertices are not adjacent, i.e., they either are in the same part or are $m$ and $m^{\prime}$. Any vertex has $\frac{l}{2}-2$ non-consecutive vertices from the other part. Then $\frac{l}{2}-2\leq 1$, so $l\in\{4,6\}$.
Figure 1: A cycle of length 6 is the sum of three cycles of length 4
In a simple cycle of length $6$ having no chords, any two opposite vertices are $m$ and $m^{\prime}$. Hence the cycle is $(m_{1},m^{\prime}_{2},m_{3},m^{\prime}_{1},m_{2},m^{\prime}_{3})$ for some pairwise distinct $m_{1},m_{2},m_{3}$. Since $n\geq 4$, there is $a\in[n]-\{m_{1},m_{2},m_{3}\}$. Now the lemma holds because (Figure 1)
$(m_{1},m_{2}^{\prime},m_{3},a^{\prime})+(m_{2},m_{3}^{\prime},m_{1},a^{\prime})+(m_{3},m_{1}^{\prime},m_{2},a^{\prime}).$
Let $t:[n]^{*2}_{\Delta}\to[n]^{*2}_{\Delta}$ be the involution switching the parts, i.e., switching $j$ and $j^{\prime}$ for every $j\in[n]$. Denote by $t$ also the induced involution on the set $H_{1}([n]^{*2}_{\Delta})$ of all $1$-cycles in $n]^{*2}_{\Delta}$.
Lemma 2.5.
The linear space $H_{1}([n]^{*2}_{\Delta})$ has a basis consisting of (the $t$-symmetric 11-cycle) $3]^{*2}_{\Delta}$ and pairs of $1$-cycles going to each other under $t$.
Proof.
Let $S$ be the set of all edges $ij^{\prime}$ of $n]^{*2}_{\Delta}$ such that $i,j>1$ and $(i,j)\neq(3,2)$. For $ij^{\prime}\in S$ let
$\widehat{ij^{\prime}}:=\begin{cases}12^{\prime}31^{\prime}ij^{\prime}&i\neq 3,j\neq 2\\ 12^{\prime}3j^{\prime}&i=3\\ 2^{\prime}31^{\prime}i&j=2\end{cases}.$
Then
$\bullet$ an edge $e^{\prime}\in S$ is contained in the cycle $\widehat{e}$ if and only if $e=e^{\prime}$;
$\bullet$ the graph $n]^{*2}_{\Delta}-S$ obtained from $n]^{*2}_{\Delta}$ by deleting all edges of $S$ is a tree.
The second of these assertions follows because the edges of $n]^{*2}_{\Delta}-S$ are exactly $32^{\prime}$ and $1i^{\prime},i1^{\prime}$, where $i>1$, so this graph is obtained by connecting with edge $32^{\prime}$ the ‘star’ trees consisting of edges $1i^{\prime}$ with $i>1$, and of edges $1^{\prime}i$ with $i>1$.
Then the cycles $\widehat{ij^{\prime}}$, $ij^{\prime}\in S$, form a basis of $H_{1}([n]^{*2}_{\Delta})$.
We have
$t\widehat{ij^{\prime}}=1i^{\prime}j1^{\prime}23^{\prime}=1i^{\prime}j1^{\prime}32^{\prime}12^{\prime}31^{\prime}23^{\prime}=\widehat{23^{\prime}}+\widehat{ji^{\prime}}.$
Then the required basis is
$3]^{*2}_{\Delta}=\widehat{23^{\prime}},\quad\widehat{ij^{\prime}},\quad t\widehat{ij^{\prime}},\quad\text{where}\quad i>j>1\quad\text{and}\quad(i,j)\neq(3,2).$
Remark 2.6.
Lemmas 2.1 and 2.2 generalize to joins of finite sets of different sizes $n_{0},\ldots,n_{k}\geq 3$. The statements generalize as follows:
$\bullet$ the joinpower $n]^{*k+1}$ is replaced with $n_{0}]*\ldots*[n_{k}]$,
$\bullet$ the cube $n]^{k+1}$ is replaced with the parallelepiped $n_{0}]\times\ldots\times[n_{k}]$,
$\bullet$ the numbers $u,v,w,a,b,c$ of parts (b) should belong to $n_{k}]$ not to $[n]$, and in the ’permuted’ relations these numbers should belong to the corresponding $n_{j}]$.
The proofs generalize by replacing
$\bullet$ $n-1]^{k+1}$ with $n_{0}-1]\times\ldots\times[n_{k}-1]$,
$\bullet$ $\{n,a_{0}\}\times\ldots\times\{n,a_{k}\}$ with $\{n_{0},a_{0}\}\times\ldots\times\{n_{k},a_{k}\}$,
$\bullet$ $(n,\ldots,n)$ with $(n_{0},\ldots,n_{k})$,
$\bullet$ the phrase ’whose $j$-th coordinate is $n$’ with the phrase ’whose $j$-th coordinate is $n_{j}$’,
$\bullet$ the phrase ’having fewer coordinates equal to $n$’ with the phrase ’having $j$-th coordinates equal to $n_{j}$ for fewer $j$’s’,
$\bullet$ $P_{j}\subset[n-1]$ with $P_{j}\subset[n_{j}-1]$,
$\bullet$ $a,b\in[n-1]$ with $a,b\in[n_{j}-1]$,
$\bullet$ $\{a,n\}$ with $\{a,n_{j}\}$,
$\bullet$ $\{b,n\}$ with $\{b,n_{j}\}$.
3 Generators in the homology of the deleted product
Of this section, only Theorems 3.1 and 3.2 are used in other sections.
In this paper $K$ is a $k$-complex (for $k=1$, a graph). Denote by $E_{k}=E_{k}(K)$ the set of all closed $k$-faces of $K$. We consider the standard cell structure on $K^{2}$, for which $E_{k}^{2}$ is the set of all closed $2k$-faces (=cells) of $K^{2}$.
A (cellular) $2k$-cycle in $K^{2}$ is a subset $C\subset E_{k}^{2}$ such that every (2k1)(2k-1)-face of $K^{2}$ is contained in an even number of $2k$-faces from $C$. In other words, such that for every (k1)(k-1)-face $a$ and $k$-face $\beta$ of $K$
$\bullet$ there is an even number of $k$-faces $\alpha$ of $K$ such that $\alpha\supset a$ and $(\alpha,\beta)\in C$, and
$\bullet$ there is an even number of $k$-faces $\alpha$ of $K$ such that $\alpha\supset a$ and $(\beta,\alpha)\in C$.
Consider the symmetry (involution) $s$ of $K^{2}$ switching the factors, and the corresponding symmetry of $2k$-cycles in $K^{2}$. A $2k$-cycle is called symmetric if $s$ preserves it. Let
$K^{\times 2}_{\Delta}:=\{(\sigma,\tau)\in E_{k}^{2}\ :\ \sigma\cap\tau=\varnothing\}\quad\text{and}\quad U_{n,k}:=([n]^{*k+1})^{\times 2}_{\Delta}$
be the combinatorial deleted product of a complex $K$, and of joinpower.
For a graph $L$ homeomorphic to $K_{5}$ or $K_{3,3}$
$\bullet$ and an edge $\sigma$ of $L$ denote by $\sigma]$ the edge of $K_{5}$ or $K_{3,3}$ ‘containing’ $\sigma$;
$\bullet$ let the economic deleted product of $L$ be $\{(\sigma,\tau)\in E_{1}(L)^{2}\ :\ [\sigma]\cap[\tau]=\varnothing\}.$
Theorem 3.1 ([Ho07, Theorem 4], cf. §5).
For a graph $K$ any symmetric $2$-cycle in $K^{\times 2}_{\Delta}$ is a sum of some of the following ones:
$\bullet$ $P\times Q+Q\times P$ for vertex-disjoint cycles $P,Q$,
$\bullet$ the economic deleted product of a subgraph homeomorphic to $K_{5}$ or to $K_{3,3}$.
Theorem 3.2.
(a) The following are $s$-symmetric $2k$-cycles in $U_{n,k}$:
$\bullet$ ‘symmetrized tori’ $P\times Q+Q\times P$ for vertex-disjoint $k$-octahedra $P,Q$;
$\bullet$ $U_{3,k}$.
(b) Any $s$-symmetric $2k$-cycle in $U_{n,k}$ is a sum of some $2k$-cycles described in (a).
Remark 3.3.
(a) For $k=1$ Theorem 3.2.b states that any symmetric $2$-cycle in $U_{n,1}=(K_{n,n})^{\times 2}_{\Delta}$ is a sum of some of the following ones: $U_{3,1}=(K_{3,3})^{\times 2}_{\Delta}$, and ‘symmetrized tori’ $P\times Q+Q\times P$ for vertex-disjoint cycles $P,Q$ of length $4$ in $K_{n,n}$. Cf. Theorem 3.1.
(b) Proof of Theorem 3.2.b provides a simpler proof (and a higher-dimensional generalization) of [FH10, the formula for $b_{2}(F(\Gamma,2))$ in Corollary 24]: $\dim H_{2k}(U_{n,k})=(n^{2}-3n+1)^{k+1}$ for $n\geq 3$.
In this paper, the following operations are considered for subspaces in vector spaces over ${\mathbb{Z}}_{2}$: linear spans ($\langle\dots\rangle$), direct sums ($\oplus$), tensor products ($\otimes$) and $\ell$-th tensor powers ($\dots{}^{\otimes\ell}$).
Cellular $j$-cycles in $K^{j}$ are defined analogously to $j=2$. Denote by $H_{j}(X)$ the group of (cellular) $j$-cycles in a (cell) complex $X$ of dimension $j$.
Theorem 3.4.
(a) There is an isomorphism
$\psi:H_{2k}(U_{n,k})\to H_{1}([n]^{*2}_{\Delta})^{\otimes k+1}\quad\text{such that}\quad\psi s=t^{\otimes k+1}\psi,\quad U_{3,k}=\psi^{-1}(([3]^{*2}_{\Delta})^{\otimes k+1}),$
and the tensor product of any $k+1$ cycles of length 4 is the $\psi$-image of the product of some two disjoint $k$-octahedra.
(b) We have
$\mathop{Ker}(I-t^{\otimes\ell})=\mathop{Im}(I+t^{\otimes\ell})\oplus\left\langle([3]^{*2}_{\Delta})^{\otimes\ell}\right\rangle\subset H_{1}([n]^{*2}_{\Delta})^{\otimes\ell}.$
Proof of Theorem 3.2.
(a) This follows because any symmetrized torus is a symmetric $2k$-cycle and by Theorem 3.4.a.
(b) The elements of $H_{1}([n]^{*2}_{\Delta})^{\otimes k+1}$ invariant under the involution $t^{\otimes k+1}$ form the subgroup $\mathop{Ker}(I-t^{\otimes k+1})$. By Lemma 2.3 $\mathop{Im}(I+t^{\otimes k+1})$ is generated by the ‘symmetrized tori’
$a_{0}\otimes\ldots\otimes a_{k}+t(a_{0})\otimes\ldots\otimes t(a_{k})\in H_{1}([n]^{*2}_{\Delta})^{\otimes k+1},$
for cycles $a_{0},\ldots,a_{k}$ of length $4$ in $n]^{*2}_{\Delta}$. Thus (b) follows by Theorem 3.4.a and Theorem 3.4.b for $\ell=k+1$. ∎
Proof of Theorem 3.4.a.
Vectors $\alpha,\beta\in[n]^{k+1}$ are said to be disjoint if $\alpha_{j}\neq\beta_{j}$ for every $j=0,\ldots,k$. For a set $X$, the set $2^{X}$ is considered with the operation of a sum modulo $2$. Let $\psi:=\varkappa|_{\ldots}\varphi|_{\ldots}$ be the composition of the isomorphisms from the following commutative diagram:
[…]
$\bullet$ $\varkappa$ is the Künneth isomorphism that maps the product $y_{0}\times\ldots\times y_{k}$ to $y_{0}\otimes\ldots\otimes y_{k}$;
$\bullet$ $\varphi$ is the isomorphism induced by the 1–1 correspondence [^8] Recall that for $2$-element subsets $P_{0},\ldots,P_{k}\subset[n]$ the $k$-faces $a_{0}*\ldots*a_{k}$, $a_{i}\in P_{i}$, of the $k$-octahedron $P_{0}*\ldots*P_{k}$ are spanned by vertices $(i,a_{i})$.
$\varphi^{\prime}:E_{2k}(U_{n,k})\to E_{k+1}(([n]^{*2}_{\Delta})^{\times k+1})\text{ given by }\varphi^{\prime}\bigl((\alpha_{0}*\ldots*\alpha_{k})\times(\beta_{0}*\ldots*\beta_{k})\bigr)=(\alpha_{0}*\beta_{0})\times\ldots\times(\alpha_{k}*\beta_{k})$
between $2k$-cells and $(k+1)$-cells for disjoint $\alpha,\beta\in[n]^{k+1}$.
If one of $\alpha_{0},\ldots,\alpha_{k},\beta_{0},\ldots,\beta_{k}$ is replaced by the empty set, then the formula for $\varphi^{\prime}$ gives a 1–1 correspondence between (2k1)(2k-1)-cells of $U_{n,k}$ and $k$-cells of $([n]^{*2}_{\Delta})^{\times k+1}$. The two 1–1 correspondences respect incidence. Thus indeed $\varphi H_{2k}(U_{n,k})=H_{k+1}(([n]^{*2}_{\Delta})^{\times k+1})$.
Then for any disjoint $\alpha,\beta\in[n]^{k+1}$, we have
$\psi\bigl((\alpha_{0}*\ldots*\alpha_{k})\times(\beta_{0}*\ldots*\beta_{k})\bigr)=(\alpha_{0}*\beta_{0})\otimes\ldots\otimes(\alpha_{k}*\beta_{k}).$ (1)
The equality $\psi s=t^{\otimes k+1}\psi$ holds because it is easily verified on generators:
$\psi s\bigl((\alpha_{0}*\ldots*\alpha_{k})\times(\beta_{0}*\ldots*\beta_{k})\bigr)=\psi\bigl((\beta_{0}*\ldots*\beta_{k})\times(\alpha_{0}*\ldots*\alpha_{k})\bigr)=$
$=(\beta_{0}*\alpha_{0})\otimes\ldots\otimes(\beta_{k}*\alpha_{k})=(t(\alpha_{0}*\beta_{0}))\otimes\ldots\otimes(t(\alpha_{k}*\beta_{k}))=$
$=t^{\otimes k+1}(\alpha_{0}*\beta_{0})\otimes\ldots\otimes(\alpha_{k}*\beta_{k})=t^{\otimes k+1}\psi\bigl((\alpha_{0}*\ldots*\alpha_{k})\times(\beta_{0}*\ldots*\beta_{k})\bigr).$
Summing up (1) over all disjoint $\alpha,\beta\in[3]^{k+1}$ gives $\psi(U_{3,k})=([3]^{*2}_{\Delta})^{\otimes k+1}$.
For any cycle $a_{j}$ of length 4 in $n]^{*2}_{\Delta}$, $j=0,\ldots,k$, there are disjoint $2$-element subsets $P_{j},Q_{j}\subset[n]$ such that $a_{j}=P_{j}*Q_{j}$. Then $P=P_{0}*\ldots*P_{k}$ and $Q=Q_{0}*\ldots*Q_{k}$ are disjoint $k$-octahedra in $n]^{*k+1}$. Summing up (1) over all $\alpha_{j}\in P_{j}$ and $\beta_{j}\in Q_{j}$ gives $\psi(P\times Q)=a_{0}\otimes\ldots\otimes a_{k}$. ∎
Proof of Theorem 3.4.b.
Take tensor $\ell$-fold products of 1-cycles of a basis from Lemma 2.5. They form a basis
$y_{-m},\ldots,y_{-1},\ y_{0}=([3]^{*2}_{\Delta})^{\otimes\ell},\ y_{1},\ldots,y_{m}\quad\text{of}\quad H_{1}([n]^{*2}_{\Delta})^{\otimes\ell}\quad\text{such that}\quad t^{\otimes\ell}y_{i}=y_{-i}.$
(I.e. the basis is involutively permuted by $t^{\otimes\ell}$, with the only fixed vector $y_{0}$.)
Thus the subgroups $\mathop{Ker}(I-t^{\otimes\ell})$ and $\mathop{Im}(I+t^{\otimes\ell})$ have bases
$y_{0},y_{1}+y_{-1},\ldots,y_{m}+y_{-m}\quad\text{and}\quad y_{1}+y_{-1},\ldots,y_{m}+y_{-m},$
respectively. This implies Theorem 3.4.b. ∎
Remark 3.5.
Theorems 3.2 and 3.4 generalize to joins of finite sets of different sizes $n_{0},\ldots,n_{k}\geq 3$.
The statements generalize by replacing (below $r=k+1,\ell$)
$\bullet$ $U_{n,k}$ with $([n_{0}]*\ldots*[n_{k}])^{\times 2}_{\Delta}$,
$\bullet$ $H_{1}([n]^{*2}_{\Delta})^{\otimes r}$ with $H_{1}([n_{0}]^{*2}_{\Delta})\otimes\ldots\otimes H_{1}([n_{r-1}]^{*2}_{\Delta})$,
$\bullet$ $t$ with the corresponding involutions $t_{j}$ of $n_{j}]^{*2}_{\Delta}$, of $2^{E_{1}([n_{j}]^{*2}_{\Delta})}$, and of $H_{1}([n_{j}]^{*2}_{\Delta})$,
$\bullet$ $t^{\otimes r}$ with $t_{0}\otimes\ldots\otimes t_{r-1}$.
The proofs generalize by replacing
$\bullet$ $n]^{k+1}$ with $n_{0}]\times\ldots\times[n_{k}]$,
$\bullet$ $([n]^{*2}_{\Delta})^{\times k+1}$ with $n_{0}]^{*2}_{\Delta}\times\ldots\times[n_{k}]^{*2}_{\Delta}$,
$\bullet$ $(2^{E_{1}([n]^{*2}_{\Delta})})^{\otimes k+1}$ with $2^{E_{1}([n_{0}]^{*2}_{\Delta})}\otimes\ldots\otimes 2^{E_{1}([n_{k}]^{*2}_{\Delta})}$,
$\bullet$ $n]^{*k+1}$ with $n_{0}]*\ldots*[n_{k}]$,
$\bullet$ the phrase ’for cycles $a_{0},\ldots,a_{k}$ of length $4$ in $n]^{*2}_{\Delta}$’ with the phrase ’for cycles $a_{0},\ldots,a_{k}$ of length $4$ in $n_{0}]^{*2}_{\Delta},\ldots,[n_{k}]^{*2}_{\Delta}$ respectively’,
$\bullet$ the phrase ’For any cycle $a_{j}$ of length 4 in $n]^{*2}_{\Delta}$, $j=0,\ldots,k$, there are disjoint $2$-element subsets $P_{j},Q_{j}\subset[n]$’ with the phrase ’For any cycle $a_{j}$ of length 4 in $n_{j}]^{*2}_{\Delta}$, $j=0,\ldots,k$, there are disjoint $2$-element subsets $P_{j},Q_{j}\subset[n_{j}]$’.
4 Proof of the ‘if’ part of Theorem 1.5
In this section $H,V$ are any linear spaces over ${\mathbb{Z}}_{2}$, and $\cdot:H\times H\to{\mathbb{Z}}_{2}$ is any symmetric bilinear form.
Lemma 4.1.
Let $P_{1},\ldots,P_{p}$ be generators of $V$, and $Y_{1},\ldots,Y_{p}\in H$ be vectors. Assume that if a sum of some $P_{j}$ is zero, then the $\cdot$-product of the sum of the corresponding $Y_{j}$, and $Y_{i}$, is zero for any $i\in[p]$.
Then there is a homomorphism $\psi:V\to H$ such that $\psi P_{a}\cdot\psi P_{b}=Y_{a}\cdot Y_{b}$ for every $a,b\in[p]$.
Proof.
 [^9] Observe that the map $P_{a}\mapsto Y_{a}$ need not extend to a homomorphism $V\to H$.
Without loss of generality, $P_{1},\ldots,P_{t}$ is a basis of $V$. Define a homomorphism $\psi:V\to H$ by setting $\psi(P_{j}):=Y_{j}$ for every $j\in[t]$.
Express any $P_{a}$ as a sum of some basic elements: $P_{a}=\sum_{j}P_{j}$. Then $\psi P_{a}-Y_{a}=\sum_{j}Y_{j}-Y_{a}$
$\bullet$ is a sum of some of vectors $Y_{1},\ldots,Y_{p}$;
$\bullet$ has zero $\cdot$-product with $Y_{i}$ for any $i\in[p]$ (by the assumption).
Hence
$\psi P_{a}\cdot\psi P_{b}-Y_{a}\cdot Y_{b}=(\psi P_{a}-Y_{a})\cdot Y_{b}+Y_{a}\cdot(\psi P_{b}-Y_{b})+(\psi P_{a}-Y_{a})\cdot(\psi P_{b}-Y_{b})=0.$
For a map $y:E_{k}\to H$ define
$\bullet$ a homomorphism $\widehat{y}:H_{k}(K)\to H$ by $\widehat{y}(A):=\sum\limits_{\sigma\in A}y(\sigma)$;
$\bullet$ $y^{2}(C,\cdot):=\sum\limits_{\{\sigma,\tau\}\in C/s}y(\sigma)\cdot y(\tau)\in{\mathbb{Z}}_{2}$, where $C$ is a symmetric $2k$-cycle in $K^{2}$.
Lemma 4.2.
(a) For any homomorphism $\psi:H_{k}(K)\to H$ there is a map $y:E_{k}\to H$ such that $\psi=\widehat{y}$.
(b) Let $y:E_{k}\to H$ be a map. For $k$-cycles $P_{1},\ldots,P_{q}$, $Q_{1},\ldots,Q_{q}$ in $K$ take a symmetric $2k$-cycle $C_{P,Q}:=\sum\limits_{j=1}^{q}(P_{j}\times Q_{j}+Q_{j}\times P_{j})$ in $K^{2}$. Then $y^{2}(C_{P,Q},\cdot)=\sum\limits_{j=1}^{q}\widehat{y}(P_{j})\cdot\widehat{y}(Q_{j})$.
Proof of (a).
(Part (a) is simple and known.) The linear operator $\psi$ extends to a linear operator $\widetilde{y}:2^{E_{k}}\to H$. Define the map $y:E_{k}\to H$ by $y(\sigma):=\widetilde{y}(\{\sigma\})$. So if $A\in H_{k}(K)$, then $\widehat{y}(A)=\sum\limits_{\sigma\in A}y(\sigma)=\sum\limits_{\sigma\in A}\widetilde{y}(\{\sigma\})=\widetilde{y}(A)=\psi(A)$. ∎
Part (b) is proved by direct calculation.
Denote by $|S|_{2}\in{\mathbb{Z}}_{2}$ the number of elements modulo 2 in a finite set $S$.
Let $f:K\to{\mathbb{R}}^{2k}$ be a general position PL map (see definition in [Sk24, Definition 1.1.6.a]). For a symmetric $2k$-cycle $C$ in $K^{\times 2}_{\Delta}$ let the $C$-van Kampen number of $K$ be
$v(C)=v_{K}(C):=\sum\limits_{\{\sigma,\tau\}\in C/s}|f\sigma\cap f\tau|_{2}\in{\mathbb{Z}}_{2},$
Recall that $v_{K}(C)$ depends only on $K,C$, not on $f$, cf. [Sk18, Remark 1.5.5.c].
[…]
Let $M$ be a closed (k1)(k-1)-connected $2k$-manifold. There is a ${\mathbb{Z}}_{2}$-embedding $K\to M$ if and only if
(R’) there is a map $y:E_{k}\to H_{k}(M)$ such that for any symmetric $2k$-cycle $C$ in $K^{\times 2}_{\Delta}$ we have $v_{K}(C)=y^{2}(C,\cap_{M})$.
Theorem 4.3 follows because (as we prove below) (R’) is equivalent to the condition (RH) of Theorem 4.4 below.
Proof of the ‘if’ part of Theorem 1.5.
Let $V=H_{k}([n]^{*k+1})$. In this paragraph we prove that there is a homomorphism $\psi:V\to H_{k}(M)$ such that
$(*)\qquad\psi P\cap_{M}\psi Q=(Y^{T}\Omega_{M}Y)_{P,Q}\quad\text{for any $k$-octahedra}\quad P,Q.$
Take $p={n\choose 2}^{k+1}$ and all different $k$-octahedra $P_{1},\ldots,P_{p}\in V$. By Lemma 2.1.a they generate $V$. Take a basis of $H_{k}(M)$ in which $\cap_{M}$ has matrix $\Omega_{M}$. For every $a\in[p]$ let $Y_{a}\in H_{k}(M)$ be the vector expressed in this basis as the column of the given matrix $Y$ corresponding to $P_{a}$. Then for any $a,b\in[p]$ we have $Y_{a}\cap_{M}Y_{b}=(Y^{T}\Omega_{M}Y)_{a,b}$. So by Lemma 2.1.b and the additivity of $Y^{T}\Omega_{M}Y$ the assumption of Lemma 4.1 is fulfilled for $H=H_{k}(M)$ and $\cdot=\cap_{M}$. This gives the required homomorphism $\psi$.
By Lemma 4.2.a there is a map $y:E_{k}\to H_{k}(M)$ such that $\psi=\widehat{y}$. Both $y^{2}(C,\cap_{M})$ and $v(C)$ are linear in $C$. So by Theorems 4.3 and 3.2.b it suffices to check the equality $y^{2}(C,\cap_{M})=v(C)$ only for $2k$-cycles $C$ from Theorem 3.2.a.
Let $C=P\times Q+Q\times P$ for vertex-disjoint $k$-octahedra $P,Q$. Then
$y^{2}(C,\cap_{M})\overset{(1)}{=}\psi(P)\cap_{M}\psi(Q)\overset{(2)}{=}0\overset{(3)}{=}|f(P)\cap f(Q)|_{2}\overset{(4)}{=}v(C),\quad\text{where}$
$\bullet$ (1) holds by Lemma 4.2.b;
$\bullet$ (2) holds by (*) and the independence;
$\bullet$ (3) holds by the well-known Parity Lemma (asserting that every two general position $k$-cycles in ${\mathbb{R}}^{2k}$ intersect at an even number of points, cf. [Sk, §1.3, §4.8 ‘Algebraic intersection number’, and Lemma 5.3.4]);
$\bullet$ (4) holds by the definition of $v(C)$.
Let $C=U_{3,k}$. Then $C=\sum(P\times Q+Q\times P)$ is the sum over all $2^{k}$ unordered pairs $\{P,Q\}$ of $k$-octahedra $P,Q\subset[3]^{*k+1}$ whose only common $k$-face is $1^{*k+1}$ [DS22, Lemma 2.3]. Hence
$y^{2}(C,\cap_{M})\overset{(1)}{=}\sum\psi(P)\cap_{M}\psi(Q)\overset{(2)}{=}S_{[3]^{*k+1},1^{*k+1}}Y^{T}\Omega_{M}Y\overset{(3)}{=}1\overset{(4)}{=}v(C),\quad\text{where}$
$\bullet$ (1) holds by Lemma 4.2.b;
$\bullet$ (2) holds by (*) and definition of $S_{X,e}$;
$\bullet$ (3) holds by the non-triviality; and
$\bullet$ (4) holds by [vK32, Satz 5] (for an alternative modern proof see [DS22, §4, proof of Lemma 4.2]). ∎
Theorem 4.4.
Let $M$ be a closed (k1)(k-1)-connected $2k$-manifold. There is a ${\mathbb{Z}}_{2}$-embedding $K\to M$ if and only if
(RH) there is a homomorphism $\psi:H_{k}(K)\to H_{k}(M)$ such that $\omega(\psi)=v_{K}$ (see the required definitions below).
This is [PT19, Proposition 21 and Theorem 15], [Sk24, Theorem 1.1.5 and Lemma 2.1.1].
Let $f:K\to{\mathbb{R}}^{2k}$ be a general position PL map. Take any pair of vertex-disjoint $k$-faces $\sigma,\tau$ of $K$. By general position the intersection $f\sigma\cap f\tau$ consists of a finite number of points. Assign to the pair $\{\sigma,\tau\}$ the residue
$\nu(f)\{\sigma,\tau\}:=|f\sigma\cap f\tau|_{2}.$
Denote by $K^{*}:=K^{\times 2}_{\Delta}/s$ the set of all unordered pairs of non-adjacent $k$-faces of $K$.
The obtained map $\nu(f):K^{*}\to{\mathbb{Z}}_{2}$ is called the (modulo 2) intersection cocycle of $f$.
We consider the standard cell structure on $K^{*}$. Let $H^{2k}(K^{*})=H^{2k}_{sym}(K^{\times 2}_{\Delta};{\mathbb{Z}}_{2})$ be the group of modulo 2 cohomology classes of maps (=cocycles) $K^{*}\to{\mathbb{Z}}_{2}$.
It is classical [vK32] that the class
$v_{K}:=[\nu(f)]\in H^{2k}(K^{*})$
is well-defined, i.e. is independent of the choice of $f$.
For a map $y:E_{k}\to H_{k}(M)$ define a map
$\omega(y):K^{*}\to{\mathbb{Z}}_{2}\quad\text{by}\quad\omega(y)\{\sigma,\tau\}:=y(\sigma)\cap_{M}y(\tau).$
By Lemma 4.2.a for any homomorphism $\psi:H_{k}(K)\to H_{k}(M)$ there is a map $y:E_{k}\to H_{k}(M)$ such that $\psi=\widehat{y}$. Define
$\omega(\psi):=[\omega(y)]\in H^{2k}(K^{*}).$
It is easy to check that $\omega(\psi)$ is well-defined, i.e. that maps $\omega(y)$ are cohomologous for two maps $y$ giving the same homomorphism $\psi=\widehat{y}$ (because such maps $y$ are cohomologous).
Remark 4.5.
We have
$\omega(\psi):=\Bigl((\psi\otimes\psi)^{*}(\cap_{M})\Bigr)|_{K^{\times 2}_{\Delta}}\in H^{2k}_{sym}(K^{\times 2}_{\Delta};{\mathbb{Z}}_{2}),\quad\text{where}$
$\bullet$ $\cap_{M}$ is considered to be an element of $\mathop{Hom}(H_{k}(M)\otimes H_{k}(M))$.
$\bullet$ we identify the following groups by the Universal Coefficients isomorphism (natural in this situation) and the Künneth isomorphism:
$\mathop{Hom}(H_{k}(K)\otimes H_{k}(K))\cong H^{k}(K)\otimes H^{k}(K)\cong H^{2k}(K\times K).$
Since $\cap_{M}$ is symmetric, $\omega(\psi)$ is indeed symmetric. See [KS21e, Remark 2.5.2b], cf. [PT19].
Denote by $\frown:H^{2k}(K^{*})\times H_{2k}(K^{*})\to{\mathbb{Z}}_{2}$ the ‘scalar’ product.
[…]
Take a general position PL map $f\colon K\to{\mathbb{R}}^{2k}$. By Lemma 4.2.a there is a map $y:E_{k}\to H_{k}(M)$ such that $\psi=\widehat{y}$. Then by (RH)
$y^{2}(C,\cap_{M})=[\omega(y)]\frown C/s=[\nu(f)]\frown C/s=v(C).$
[…]
Take a general position PL map $f\colon K\to{\mathbb{R}}^{2k}$. Let $\psi=\widehat{y}$. Then
$\omega(\psi)=[\omega(y)]=[\nu(f)]=v_{K}.$
Here the middle equality holds because the ‘scalar’ product $\frown$ is nondegenerate, and for any symmetric $2k$-cycle $C$ in $K^{\times 2}_{\Delta}$ by (R’) we have
$\sum_{\{\sigma,\tau\}\in C/s}(\omega(y)-\nu(f))\{\sigma,\tau\}=y^{2}(C,\cap_{M})-v(C)=0.$
Remark 4.6.
Theorem 1.5 generalizes to joins of finite sets of different sizes $n_{0},\ldots,n_{k}\geq 3$ (see Remark 1.6.c). The proof of the ‘if’ part of Theorem 1.5 generalizes by replacing $n]^{*k+1}$ with $n_{0}]*\ldots*[n_{k}]$ and $p={n\choose 2}^{k+1}$’ with $p={n_{0}\choose 2}\ldots{n_{k}\choose 2}$’.
5 On the paper [Sa91]
In [Sa91] proof of [Sa91, 3.4.1] (Theorem 3.1) has the following two gaps, see Remarks 5.1.a and 5.2.b. We use the notation from [Sa91]: $K_{*}$ is the deleted product ($K^{\times 2}_{\Delta}$ of this paper).
Remark 5.1.
For edges $\alpha,\beta$ of $K$ denote by $K_{\alpha,\beta}$ the graph obtained from $K$ by subdividing both edges $\alpha,\beta$, and then identifying both appearing points of degree 2 to a new vertex denoted by 0.
(a) In [Sa91, bottom of p. 86], the statement
(*) ’But $L_{*}$ has less symmetric circuits than $(K_{z})_{*}$
is not proved. Here
(0) ‘symmetric circuit’ is non-standard terminology for ‘symmetric (cellular) 2-cycle’;
(1) $K_{z}$ is a graph that is the support of a minimal symmetric 2-cycle $z$ in $(K_{z})_{*}$, and $(K_{z})_{*}$ contains a minimal symmetric 2-cycle $w\neq z$;
(2) $\alpha\times\beta$ is a cell from wzw-z, and $L:=(K_{z})_{\alpha,\beta}$.
A ‘trivial proof’ of (*) fails by (b). Although (b) is not a counterexample to (*), (b) shows that the proof of (*) cannot be trivial because it has to use the assumptions (1) and (2).
(b) There is a graph $K$ and its disjoint edges $\alpha,\beta$ such that $K_{\alpha,\beta,*}$ contains a symmetric 2-cycle not contained in $K_{*}$.
Indeed, let $K$ be the union of the cycle $12345$ and chord $14$. In $K_{12,34,*}$ the symmetric 2-cycle $C:=023\times 145+145\times 023$ contains $02\times 14$. The cell $02\times 14$ is not contained in $K_{*}$. Therefore, $C$ is not contained in $K_{*}$. [^10] Alternatively, let $K$ be the union of $K_{4}$ and the paths $152$ and $364$. Then $K$ is planar but $K_{12,34}$ is homeomorphic to $K_{5}$, so $K_{12,34}$ contains a symmetric 2-cycle not contained in $K_{*}$.
(c) In [SSv3, §5, proof of Lemma 5.2], there is the following gap similar to the one indicated in (a). The following statement is not proved (and by (b) is incorrect without quite technical assumptions on the graph $L$ and the edges $e_{1},e_{2}$ for which $L_{0}=L_{e_{1},e_{2}}$):
… symmetric $2$-cycles in $(L_{0})^{\times 2}_{\Delta}$ are exactly symmetric $2$-cycles in $L^{\times 2}_{\Delta}$ not containing $(e_{1},e_{2})$.
Remark 5.2.
(a) Lemma. Define a homomorphism [^11] The argument [Sa91, 3.3.1 and its proof] only uses that $v$ is a homomorphism trivial on symmetrized tori $c_{1}\times c_{2}+c_{2}\times c_{1}$. However, that argument has a gap, see (b). $v:H_{2}^{s}(K_{*};{\mathbb{Z}}_{2})\to{\mathbb{Z}}_{2}$ by $v(c):=\left<o(K),c\right$>. If $v$ is non-zero, then there is a subgraph $G$ of $K$ such that $G$ is homeomorphic either to $K_{5}$ or to $K_{3,3}$, and the restriction of $v$ to $H_{2}^{s}(G_{*};{\mathbb{Z}}_{2})$ is non-zero.
This lemma holds by Kuratowski criterion. Observe that [Sa91, 3.3.1 and its proof] intends to reprove Kuratowski criterion.
(b) ‘…[The following] results from the above discussion’ before [Sa91, 3.4.1] is unjustified. Indeed, ‘the above discussion’ [Sa91, 3.3.1 and its proof] amounts to Lemma (a). In order to conclude [Sa91, 3.4.1] from Lemma (a), one needs (an argument relating symmetric and ordinary homology, and) the following statement [Sa91, after 3.4.2]:
The toral circuits generate the kernel of the linear functional $\left<o(K),\cdot\right>:H_{2}^{s}(K_{*};{\mathbb{Z}}_{2})\to{\mathbb{Z}}_{2}$.
This statement is non-trivial, but is presented in [Sa91, after 3.4.2] without proof or reference. (The statement is correct by [Ho07] which proves [Sa91, 3.4.1] itself. See also Remark 5.3.a.)
An attempt to recover this gap is presented in [SSv3, §5], by turning the argument [Sa91, 3.3.1 and its proof] (essentially attempting to prove Lemma (a) independently of Kuratowski criterion) to the proof of (the symmetrized version of) [Sa91, 3.4.1] itself. Unfortunately, that attempt reproduced the gap 5.1.(a) as described in 5.1.(c).
Remark 5.3.
(a) The following corollary of [Sa91, 3.4.2] is wrong (even for connected graphs):
For any graph $K$ there is a cellular 2-cycle $C$ in $K_{*}$ such that any cellular 2-cycle in $K_{*}$ is a sum of some of the following cellular 2-cycles: $C$ and products of vertex-disjoint cycles.
Indeed, take disjoint union of two copies of $K_{5}$. (Make the disjoint union connected by joining the copies with an edge.) Cf. [HNT].
(b) For a description and recovery of a gap in [Sa91, §2] see [Sc13, footnote 6] and [MPS, proof of Theorem 3.1].
6 Appendix: some alternative proofs
A direct proof of Lemma 2.1.b for $k=1$.
Denote by $\alpha(a,b;u,v,w)$ and $\beta(a,b,c;u,v)$ the relations from the statement. For any $b,u,v\in[n]-\{1\}$ such that $u\neq v$, one has
$\text{for any }a\in[n]-\{b\}$ $au^{\prime}bv^{\prime}=a1^{\prime}bu^{\prime}+a1^{\prime}bv^{\prime}+\alpha(a,b;1,u,v);$
[…]
Hence any linear relation on cycles of length 4 in $K_{n,n}$ can be reduced (by adding some of $\alpha$’s and $\beta$’s) to a relation including only cycles $11^{\prime}bv^{\prime}$. The latter relation is trivial because for any $a,b,u,v\in[n]-\{1\}$ the edge $\{a,u^{\prime}\}$ belongs to the cycle $11^{\prime}bv^{\prime}$ if and only if $a=b$ and $u=v$. ∎
Alternative proof of Theorem 3.4.b.
Induction on $\ell$.
[…]
In the following paragraph we prove that any $1$-cycle in $K_{n}$ is a sum of some cycles of length $4$ and, possibly, of the cycle $(1,2,3)$.
The sum of two cycles of length $3$ having a common edge is a cycle of length $4$. So if a cycle $\alpha$ of length $3$ is a sum of $(1,2,3)$ and cycles of length 4, then any cycle of length $3$ having a common edge with $\alpha$ is such. Hence all cycles of length $3$ are such. Now the required assertion follows because by Statement 2.4 any $1$-cycle in $K_{n}$ is a sum of some cycles of length 3.
[…]
The inductive step $\ell\to\ell+1$ is obtained by applying the following Lemma 6.1 to
$W_{1}=H_{1}(\widetilde{K_{n}}),\quad W_{2}=H_{1}(\widetilde{K_{n}})^{\otimes\ell},\quad x_{1}=\widetilde{K_{3}},\quad x_{2}=\widetilde{K_{3}}^{\otimes\ell},\quad A_{1}=t,\quad A_{2}=t^{\otimes\ell}.$
Lemma 6.1.
For $j=1,2$ let $W_{j}$ be a vector space over ${\mathbb{Z}}_{2}$, $x_{j}\in W_{j}-\{0\}$ a vector, and $A_{j}$ an operator in $W_{j}$ such that $\mathop{Ker}(I+A_{j})=\mathop{Im}(I+A_{j})\oplus\left\langle x_{j}\right\rangle$. Then the operator $A:=A_{1}\otimes A_{2}$ in $W:=W_{1}\otimes W_{2}$ satisfies $\mathop{Ker}(I+A)=\mathop{Im}(I+A)\oplus\left\langle x_{1}\otimes x_{2}\right\rangle$.
Proof.
Let $N_{j}:=I+A_{j}$. Then $\mathop{Ker}N_{j}=(\mathop{Im}N_{j})\oplus\left\langle x_{j}\right\rangle$. Hence $N_{j}^{2}=0$.
Denote by $J_{m}(\lambda)$ the Jordan cell of size $m$ with eigenvalue $\lambda$. Take the Jordan form of $N_{j}$, i.e., represent $W_{j}$ as a direct sum of $N_{j}$-invariant subspaces $W_{j,p}$, $0\leq p\leq m_{j}$, such that
$\bullet$ $W_{j,0}=\left\langle x_{j}\right\rangle$, $N_{j}W_{j,0}=0$;
$\bullet$ for every $p>0$, the subspace $W_{j,p}$ is two-dimensional, and the operator $N_{j}$ acts on it with matrix $J_{2}(0)$ (in some basis).
Hence the operator $A_{j}$ also preserves every subspace $W_{j,p}$, acting on it identically if $p=0$, and with matrix $J_{2}(1)$ (in some basis) otherwise. So the space $W$ is a direct sum of $A$-invariant subspaces $W_{p_{1},p_{2}}:=W_{1,p_{1}}\otimes W_{2,p_{2}}$, $0\leq p_{j}\leq m_{j}$. Now the lemma holds because $W_{0,0}=\left\langle x_{1}\otimes x_{2}\right\rangle$, $(I+A)W_{0,0}=0$, and
(*)   $\mathop{Ker}(I+A|_{U})=\mathop{Im}(I+A|_{U})$ on every other summand $U:=W_{p_{1},p_{2}}$.
Let us prove assertion (*). Denote by $M(X)$ the matrix of an operator $X$ in the considered basis.
If either $p_{1}>0=p_{2}$ or $p_{2}>0=p_{1}$, then $U$ has a basis $e_{1},e_{2}$ in which $M(A|_{U})=J_{2}(1)$. So $M(I+A|_{U})=J_{2}(0)$. Thus, $\mathop{Ker}(I+A|_{U})=\mathop{Im}(I+A|_{U})=\left\langle e_{1}\right\rangle$.
If $p_{1},p_{2}>0$, then $U$ has a basis $e_{1},e_{2},e_{3},e_{4}$ in which
$M(A|_{U})=\begin{pmatrix}J_{2}(1)&J_{2}(1)\\ 0&J_{2}(1)\end{pmatrix}=\begin{pmatrix}1&1&1&1\\ 0&1&0&1\\ 0&0&1&1\\ 0&0&0&1\end{pmatrix},\quad\text{so}\quad M(I+A|_{U})=\begin{pmatrix}0&1&1&1\\ 0&0&0&1\\ 0&0&0&1\\ 0&0&0&0\end{pmatrix}$.
Hence $\mathop{Ker}(I+A|_{U})=\mathop{Im}(I+A|_{U})=\left\langle e_{1},e_{2}+e_{3}\right\rangle$. ∎
Remark 6.2.
Denote by $A^{\prime}$ a copy of $A$. The simplicial deleted join $G^{*2}_{\Delta}$ of a complex $G=(V,F)$ is the complex with the set $V\sqcup V^{\prime}$ of vertices and the set $\{\sigma*\tau^{\prime}\ :\ \sigma,\tau\in F,\ \sigma\cap\tau=\varnothing\}$ of faces (here $\sigma$ and $\tau$ are considered as faces of a simplex, so that $\sigma*\tau^{\prime}$ is the face spanned by all the vertices of $\sigma$ and of $\tau^{\prime}$; if $\sigma$ and $\tau$ are considered as subsets, then $\sigma*\tau^{\prime}$ should be replaced by $\sigma\sqcup\tau^{\prime}$).
Analogously to the isomorphism $\varphi$ from the proof of Theorem 3.4.a one proves that for complexes $A,B$ of dimensions $a,b$ the map $(A*B)^{\times 2}_{\Delta}\to A^{*2}_{\Delta}\times B^{*2}_{\Delta}$ given by $(\alpha_{1}*\beta_{1})\times(\alpha_{2}*\beta_{2})\mapsto(\alpha_{1}*\alpha_{2})\times(\beta_{1}*\beta_{2})$ on $(a+b+2)$-cells induces an isomorphism of the groups $H_{a+b+2}$ of cellular $(a+b+2)$-cycles. This is the analogue of [Ma03, Lemma 5.5.2] for deleted products.
An alternative definition of the isomorphism $\varphi$ is the composition
$H_{2}((K_{n,n})^{\times 2}_{\Delta})\overset{(a)}{\cong}H_{3}((K_{n,n})^{*2}_{\Delta})\overset{(b)}{\cong}H_{3}(\widetilde{K_{n}}*\widetilde{K_{n}})\overset{(c)}{\cong}H_{2}(\widetilde{K_{n}}\times\widetilde{K_{n}}),\quad\text{where}$
$\bullet$ isomorphism (a) holds by [PS20, Lemma 4d], cf. [Sk, 7.2.8c and 7.2.3b]; it is induced by the map $\sigma\times\tau\mapsto\sigma*\tau$ from the set of $2$-cells to the set of $3$-faces;
$\bullet$ isomorphism (b) holds because $K_{n,n}=[n]^{*2}$ and $\widetilde{K_{n}}=[n]^{*2}_{\Delta}$, so $(K_{n,n})^{*2}_{\Delta}\cong\widetilde{K_{n}}*\widetilde{K_{n}}$ [Ma03, Lemma 5.5.2]; the latter homeomorphism maps $(\alpha_{1}*\alpha_{2})*(\beta_{1}*\beta_{2})$ to $(\alpha_{1}*\beta_{1})*(\alpha_{2}*\beta_{2})$ for $\alpha,\beta\in[n]^{2}$ such that $\alpha_{j}\neq\beta_{j}$ for every $j\in[2]$;
$\bullet$ isomorphism (c) is induced by the map $\sigma*\tau\mapsto\sigma\times\tau$ from the set of $3$-faces to the set of $2$-cells.